### v1

# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

Abstract
Approaching the challenge to understand fundamental physical constants (FPCs) started with understanding the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss first and second-order phase transitions and emerging new states including long-lived metastable states. We also discuss transitions between metastable states, chemical reactions and their products. These examples show that different values of FPCs would disable these transitions, and new phases as a result of these transitions would not emerge. This suggests that the current values of FPCs, by enabling transitions which give rise to new states, act to increase the phase space and entropy.
1 Introduction
Fundamental physical constants (FPCs), determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Uzan (2003, 2011); Tiesinga et al. (2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
Figure 1: Different scales of structure in the physical world including the range of typical sizes studied in condensed matter physics. Fairly recently, fundamental physical constants were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli and velocity gradients that can be set up in cells using biochemical energy (see discussion in the Introduction). In section 3, we discuss a wider role of fundamental constants affecting observability and operation of entire phenomena and effects, including second-order structural and superconducting transitions, first-order melting transition, metastable transitions and chemical and other reactions.
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006). The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983). We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019)? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
The top in Figure 1 involves cosmological length scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019). The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
Compared to the above high-energy processes, the discussion of the role of FPCs at the scale of condensed matter has been scarce - this is the scale spanning 108-109 orders of magnitude in Figure 1 or about 20% of the logarithmic scale between the nuclei and the size of Universe. This has started to change fairly recently.
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b). This has deepened our understanding of these properties and their physical models.
In this Perspective, we discuss a wider role of FPCs in condensed matter: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects. In the next section, we discuss why discussing FPCs at the condensed matter scale is important. This is followed by several examples in condensed matter physics showing how FPCs set observability and operation of phase transitions, the existence of different phases and states of matter emerging as a result of these transitions, including numerous metastable states. We also discuss chemical reactions and their products. These examples show that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This suggests that the current values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy. This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
2 Why filling the condensed matter gap is important
There are several reasons why understanding the role of FPCs at condensed matter length and energy scales and filling the condensed matter gap in Figure 1 is important. First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs Adams (2019). On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
Third, understanding the role of FPCs in condensed matter physics is an important enterprise on its own. It enriches the field itself and deepens understanding condensed matter phenomena at the most basic level. It also provides links and relationships to other areas of physics.
The fourth and final reason discussed in this Perspective is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales. This can lead to new insights related to the origin of FPCs and their values.
In the next section, we discuss several examples where FPCs set observability and operation of condensed matter effects and phenomena.
3 Observability of phase transitions and existence of different states
3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first is the set of FPCs. The second type are the environmental parameters such as, for example, temperature and pressure at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the temperature of the planet. In several examples in this section, we consider a certain range of environmental parameters (for example, the temperature range 102-103 K at typical planetary conditions in the Solar system) and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
We will be using and discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
$\beta=\frac{m_{p}}{m_{e}}$ (1)
and the fine-structure constant
$\alpha=\frac{e^{2}}{\hbar c}$ (2)
where $c$ is the speed of light in vacuum.
$\alpha$ and $\beta$ are of particular importance when discussing processes at different length scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter states, solids and liquids, are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
$a_{\rm B}=\frac{4\pi\epsilon_{0}\hbar^{2}}{m_{e}e^{2}}$ (3)
and
$E_{\rm R}=\frac{m_{e}e^{4}}{32\pi^{2}\epsilon_{0}^{2}\hbar^{2}}$ (4)
$a_{\rm B}$ is about 0.5 Å and gives a shorter range of typical interatomic separations condensed phases in the range 1-3 Å. $E_{\rm R}$ is about 13.6 eV and gives a larger value of cohesive energy in condensed phases in the typical range 1-10 eV.
3.2 Second-order phase transitions
Our first case study are the second-order phase structural transitions in solids. We consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. For simplicity, we consider two phases only, quartz and $\beta$-cristobalite, its high-temperature (around 1800 K) phase Hammonds et al. (1996); Dove (1997).
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
$S_{s}(T\gg\hbar\omega_{\rm D})=3N\left(\ln\frac{T}{\hbar\bar{\omega}}+1\right)$ (5)
where $\bar{\omega}$ is the average phonon frequency defined as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies and $N$ is the number of atoms. Here and below, $k_{\rm B}=1$.
Compared to quartz, cristobalite has a more symmetric and open structure and a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy (5) as a result.
The entropy (5) reflects the phase space available to $3N$ phonons in solids, and the increase of entropy of cristobalite is due to the phonon softening. However, the same entropy increase can be seen directly as the increase of the phase space of particle coordinates $q$ and momenta $p$, $\Delta\Gamma$. $\Delta\Gamma$ is:
$\Delta\Gamma=\frac{1}{(2\pi\hbar)^{3N}}\prod\limits_{i}\Delta p_{i}\Delta q_{i}$ (6)
Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq. (6) and taking the relation between $\Delta\Gamma$ and $S$ as
$S=\ln\Delta\Gamma$ (7)
gives the entropy term $S\propto 3N\ln\frac{T}{\hbar\bar{\omega}}$ as in Eq. (5).
We see that smaller $\omega_{i}$ due to softer phonons in cristobalite increase $\langle\Delta p_{i}\rangle^{2}$ and $\langle\Delta q_{i}\rangle^{2}$. This increases $\Delta\Gamma$ and $S$ of cristobalite compared to quartz.
At room temperature or any other temperature below the phase transition, cristobalite is a long-lived metastable phase of silica. Lets compare the entropy of quartz with the number of atoms $N$, $S_{Q}$, and the entropy of quartz-cristobalite mixture, $S_{Q-C}$, consisting of any proportions of the two phases and having the same total number of atoms $N$. Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined and used in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system (or system sizes small enough), the system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence, at room or other fixed temperature where both phases exist, we have
$S_{Q-C}>S_{Q}$ (8)
We therefore find that the increase of the phase space and entropy (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase. Without a phase transition operating in the considered range, the entropy will increase according to Eq. (5) within the single quartz phase but without the frequency softening and the associated increase of the phase space (we do not consider softening of phonon frequencies with temperature due to usual anharnomicity).
We now observe that if we consider different values of FPCs, the phase transition would not operate. We recall that the upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants ($\omega_{m}$ is sometimes referred to as the “rattling” frequency) Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ as
$\frac{\hbar\omega_{m}}{E}=\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (9)
where $m$ is the atom mass.
Setting $E=E_{\rm R}$ gives
$\omega_{m}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (10)
which has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
$\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (11)
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (10)-(11). Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain the rest of essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
Larger $\omega_{m}$ implies larger $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) is no longer applicable because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is Landau and Lifshitz (1970)
$S_{s}(T\ll\hbar\omega_{\rm D})=\frac{4\pi^{4}}{5}N\left(\frac{T}{\hbar\omega_{\rm D}}\right)^{3}$ (12)
$S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$. This means that the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition. The phase transition becomes inoperative.
Therefore, the current values of fundamental constants, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy in Eq. (8).
Figure 2 illustrates this point for solid minerals including quartz considered in this case study. A binary quartz-cristobalite system considered with increased entropy (8) is readily generalised to a collection of many systems which exist in long-lived metastable states (see Section 3.5 for a detailed discussion of metastable states). Note that softening and the associated increase of the phase space in high-temperature phases is generally expected because these phases tend to be more symmetric and have a larger number of soft phonons as a result Hammonds et al. (1996); Dove (1997).
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. The mixture of quartz and metastable cristobalite at the same temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms). Phase transitions between different phases enable collections of systems with increased phase space and entropy at the same temperature. Shown minerals are for illustration only.
This case study gives us several insights. First, FPCs govern important effects in condensed matter and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, can increase the phase space and entropy.
3.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked Trachenko et al. whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
In the Migdal-Eliashberg theory, the upper limit to $T_{c}$ can be written as $T_{c}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (11) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum at a certain $\lambda$ because $\lambda$ larger typically above 1-2 result in instabilities and associated lowering of $T_{c}$. Using $C$ in the range 0.1-0.2 indicated by experiments, theory and modelling and Eq. (11) gives the upper bound on $T_{c}$, $T_{c}^{u}$, as Trachenko et al.
$T_{c}^{u}=C\omega_{m}^{u}=C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\approx 10^{2}-10^{3}\,\rm{K}$ (13)
Using Eq. (13), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved at our environmental conditions.
Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature. Consistent with Eq. (13) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above. We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
3.4 First-order phase transitions: solids, liquids, gases
Three basic states of matter, solids, liquids and gases, are represented by a phase diagram with three first-order phase transition lines as shown in Figure 3. In this section, we discuss how the observability of different states of matter and transitions between those states at our environmental conditions is related to FPCs.
Figure 3: Three states of matter, solids, liquids and gases, separated by three phase transition lines. The triple and critical points are shown.
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) set by FPCs in Eq. (4) because microscopically, melting involves breaking chemical bonds. Experimentally, melting temperatures correlate with cohesive energies and are on the order of $\omega_{\rm D}$ in the range 102-103 K for many systems Ubbelohde (1978); Grimvall and Sjödin (1974). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs in Eq. (10). This then gives the following role of FPCs in setting the states of matter at environmental conditions.
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only. As in the case of the second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
The overlap of our environmental conditions of pressure and temperature with melting and boiling temperatures set by FPCs gives rise to two states important for life: at our environmental conditions, H2O can exist as water and carbon-based substances exist as bonded solidlike structures. Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only at environmental conditions and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures substances would not exist as cohesive states but as the gas state only.
We note that classical entropies of solids and gases contain FPCs explicitly. Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq. (10) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
$S_{s}\propto\ln\frac{\hbar^{2}m_{p}^{\frac{1}{2}}}{e^{4}m_{e}^{\frac{3}{2}}}$ (14)
The entropy of the classical gas is Landau and Lifshitz (1970)
$S_{g}=\frac{3}{2}N\left(\ln T+1\right)+N\left(\ln v+\ln\left(\frac{m}{2\pi\hbar^{2}}\right)^{\frac{3}{2}}+1\right)$ (15)
where $v$ is the volume per atom and $m$ is the atom mass.
With $v\propto a_{\rm B}^{3}$ and $m\propto m_{p}$,
$S_{g}\propto\ln\frac{\hbar^{3}m_{p}^{\frac{3}{2}}}{e^{6}m_{e}^{3}}$ (16)
$S_{g}$ in Eq. (16) is more sensitive to the variation of FPCs as compared to $S_{s}$ in Eq. (14) so that larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster (note that larger $\hbar$ and smaller $m_{e}$ and $e$ decrease the cohesive energy in Eq. (4), favouring gases over solids). Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective, however this may change if the same solid can exist as a mixture of the stable and many long-lived metastable states (see Section 3.2 and next section for a more detailed discussion).
Differently to solid and gas entropies we considered, no closed equation for the liquid entropy exists which corresponds to experimental properties of real liquids such as $c_{v}=3$ close to melting ($c_{v}$ is the constant-volume specific heat). This is due to the fundamental no-small parameter problem of the liquid theory (see Ref. Trachenko (2024) for a quick review and Refs. Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion). Existing models such hard-sphere and Van der Waals models and their extensions give $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically. However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: energy of the system is the sum of energies of its excitations Lifshitz and Pitaevskii (2006). In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024). The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy. Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.
3.5 Metastable states and chemical reactions
In three previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions and phases assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the minimum of the free energy. Nevertheless, metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition point corresponds to the equality of the free energies of both phases (see Ref. Brazhkin (2024) for review). When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2024).
As in the case of quartz-cristobalite mixture discussed in Section 3.2, the ability of a given substance to exist in many long-lived metastable states increases the phase space and entropy. Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
Transitions between metastable states involves the same processes as transitions between equilibrium states. These processes are governed by the equality of the free energy of long-lived metastable states at the transition point. Hence, similarly to what we find for equilibrium transitions in earlier sections, the observability of transitions betweeen different metastable states and hence the existence of numerous metastable phases is set by FPCs. Similarly to what we saw in Section 3.2, the phase space and entropy at a given temperature increase due to many long-lived metastable states emerging as a result of these transitions.
The same assertion of observability of transitions and different phases applies to chemical reactions and reactions products. Chemical reactions involve changes of chemical bonding whose characteristic energies are set by FPCs. Hence these reactions are governed by the interplay of characteristic cohesive energies set by FPCs and environmental parameters. For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be unobservable at given environmental parameters.
4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and symmetry Zee (2007). FPCs would then be calculable from this theory.
In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the phase space and entropy. Metastable states were important in this discussion because at given environmental parameters (e.g., temperature), the collection of many long-lived metastable states (as opposed to just one thermodynamically stable phase) give the increase of the phase space and entropy. We discussed numerous metastable states in condensed matter physics and chemistry in Section 3. However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales of matter in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the current age of the Universe Brazhkin (2024)). These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading new states such as the nuclear synthesis in stars).
We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and long-lived metastable phases emerging, with the associated increase of the phase space and entropy. We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy. This suggests that the current values of fundamental constants act to increase the entropy. This insight is general and can be applied to different levels of structure in the Universe in Figure 1. This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism (see Ref. Trachenko (2023d) for the review of related earlier discussions) where entropy increase is analogous to selection due to environmental pressure in biological evolution. The mechanism of variability (fluctuations) of FPCs, analogous to genetic mutation in biological evolution, is currently unknown, although several mechanisms were discussed (see, e.g., Refs. Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013)).
Cyclic or oscillatory models of the Universe (see, e.g., Refs. Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref. Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle. In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates. The size of the oscillating Universe in Ref. Trachenko (2017) at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space or close to it. Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g. more baryons and stars, more efficient nuclear synthesis in stars and so on). The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs. Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space. Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles. There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (this selection applies to FPCs and FPC-enabled structures existing at the time). This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence. This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact. This invites an input from the optimisation theory.
One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy. If most of the current entropy is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
$S_{BH}=4\pi\frac{G}{\hbar c}M_{BH}^{2}$ (17)
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion of smaller BHs).
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe, identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and refining our models setting constraints on FPCs Adams (2019). Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
5 Discussion and summary
Approaching the challenge to explain fundamental constants started with understanding the role these constants play. This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties. This has enriched condensed matter physics and deepened our understanding of its key phenomena.
Here, we discussed a wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena. We gave examples of operation of first- and second-order phase transitions and emerging new phases, existence of numerous metastable phases, chemical reactions, their products and so on. We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge. This implies that the current values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
This list of effects and phenomena in condensed matter can be probably extended significantly. The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail. Physical models of condensed matter have been tested and cross-checked in numerous different experiments and simulations. In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref. Adams (2019) for a review of these challenges).
If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe. The insight that FPCs act to increase the entropy is one such example. This and similar insights can provide useful links between different areas of physics and open up new discussions.
Last by not least, deeper understanding the role of FPCs in condensed matter physics is important on its own. Understanding lower and upper bounds of different condensed matter properties including viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) has deepened our understanding of how these properties operate and benefited our physical models of these properties. Understanding how FPCs govern the observability and operation of entirely new effect and phenomena is another useful insight enhancing our understanding of condensed matter physics.
I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.

### v2

@@
 Abstract
-Approaching the challenge to understand fundamental physical constants (FPCs) started with understanding the role these constants play in high-energy nuclear physics and astrophysics.
+Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics.
 Condensed matter physics was relatively unexplored in this regard.
@@
 (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al.
-(2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b).
+(2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b) and superconducting temperature Trachenko et al.
+.
 This has deepened our understanding of these properties and their physical models.
@@
 This can lead to new insights related to the origin of FPCs and their values.
-In the next section, we discuss several examples where FPCs set observability and operation of condensed matter effects and phenomena.
-3 Observability of phase transitions and existence of different states
+In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
+3 Observability and operation of phase and other transitions. Existence of different phases and states
 3.1 Fundamental and environmental parameters
@@
 Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions.
-For simplicity, we consider two phases only, quartz and $\beta$-cristobalite, its high-temperature (around 1800 K) phase Hammonds et al.
+Let’s first consider $\beta$-cristobalite, the high-temperature phase emerging at the phase transition around 1800 K Hammonds et al.
 (1996); Dove (1997).
@@
 Here and below, $k_{\rm B}=1$.
-Compared to quartz, cristobalite has a more symmetric and open structure and a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al.
+As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz.
+As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al.
 (1996); Dove (1997).
-This gives smaller $\bar{\omega}$ in cristobalite and larger entropy (5) as a result.
+This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq.
+(5).
 The entropy (5) reflects the phase space available to $3N$ phonons in solids, and the increase of entropy of cristobalite is due to the phonon softening.
@@
 This increases $\Delta\Gamma$ and $S$ of cristobalite compared to quartz.
-At room temperature or any other temperature below the phase transition, cristobalite is a long-lived metastable phase of silica.
-Lets compare the entropy of quartz with the number of atoms $N$, $S_{Q}$, and the entropy of quartz-cristobalite mixture, $S_{Q-C}$, consisting of any proportions of the two phases and having the same total number of atoms $N$.
+Cristobalite is a long-lived metastable phase of silica which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al.
+(1996); Dove (1997).
+According to Eq.
+(5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity).
 Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems.
-These functions are defined and used in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system (or system sizes small enough), the system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970).
+These functions are defined in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970).
 This is discussed in Section 3.5 in more detail.
-$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence, at room or other fixed temperature where both phases exist, we have
+As a more general case compared to pure metastable cristobalite, let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below the transition point where both phases co-exist, $S_{Q-C}$.
+This mixture consists of any proportion of the two phases with the total number of atoms $N$.
+We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$.
+$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier.
+Hence,
 $S_{Q-C}>S_{Q}$ (8)
-We therefore find that the increase of the phase space and entropy (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase.
-Without a phase transition operating in the considered range, the entropy will increase according to Eq.
-(5) within the single quartz phase but without the frequency softening and the associated increase of the phase space (we do not consider softening of phonon frequencies with temperature due to usual anharnomicity).
+We therefore find that the entropy increase (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase.
+The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state at low $T$.
+Without a phase transition operating in the considered range, the entropy would increase according to Eq.
+(5) within the single quartz phase but without the higher entropy contribution from the second metastable phase.
 We now observe that if we consider different values of FPCs, the phase transition would not operate.
-We recall that the upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants ($\omega_{m}$ is sometimes referred to as the “rattling” frequency) Trachenko et al.
+The upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants Trachenko et al.
 .
@@
 $\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (11)
+$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1).
+As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
 Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs.
@@
 Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs.
-(1) and (2) unchanged and hence maintain the rest of essential processes (e.g.
+(1) and (2) unchanged and hence maintain all preceding essential processes (e.g.
 those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
@@
 Therefore, the current values of fundamental constants, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy in Eq. (8).
-Figure 2 illustrates this point for solid minerals including quartz considered in this case study.
-A binary quartz-cristobalite system considered with increased entropy (8) is readily generalised to a collection of many systems which exist in long-lived metastable states (see Section 3.5 for a detailed discussion of metastable states).
-Note that softening and the associated increase of the phase space in high-temperature phases is generally expected because these phases tend to be more symmetric and have a larger number of soft phonons as a result Hammonds et al.
-(1996); Dove (1997).
+A binary quartz-cristobalite mixture with increased entropy (8) is readily generalised to a collection of many systems existing in long-lived metastable states.
+Figure 2 illustrates this point.
+The increase of the phase space and entropy in these metastable states is generally expected because (a) high-temperature phases tend to be more symmetric Landau and Lifshitz (1970) with more open structure and have a larger number of soft phonons as a result Hammonds et al.
+(1996); Dove (1997) and (b) the increased phase space and entropy in the high-temperature phases become “frozen-in” in metastable states at low $T$.
 Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition.
-The mixture of quartz and metastable cristobalite at the same temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms).
-Phase transitions between different phases enable collections of systems with increased phase space and entropy at the same temperature.
+The mixture of quartz and metastable cristobalite at low temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms).
+Phase transitions between different states enable sets of long-lived metastable systems with higher combined entropy at the same temperature.
 Shown minerals are for illustration only.
 This case study gives us several insights.
-First, FPCs govern important effects in condensed matter and define the regime in which the system is, classical or quantum.
+First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum.
 Second, FPCs affect the observability and operation of entire effects such as phase transitions.
@@
 Fundamental limit here specifically means the limit in terms of FPCs.
-In the Migdal-Eliashberg theory, the upper limit to $T_{c}$ can be written as $T_{c}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq.
+Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq.
 (11) Trachenko et al.
 .
-$C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum at a certain $\lambda$ because $\lambda$ larger typically above 1-2 result in instabilities and associated lowering of $T_{c}$.
-Using $C$ in the range 0.1-0.2 indicated by experiments, theory and modelling and Eq.
-(11) gives the upper bound on $T_{c}$, $T_{c}^{u}$, as Trachenko et al.
+$C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$.
+This range of $\lambda$ corresponds to $C=0.1-0.2$.
+Using Eq.
+(11), this gives $T_{c}^{u}$ as Trachenko et al.
 $T_{c}^{u}=C\omega_{m}^{u}=C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\approx 10^{2}-10^{3}\,\rm{K}$ (13)
@@
 (13), we ask how would the observability of superconductivity change were fundamental constants to take different values?
-Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved at our environmental conditions.
+Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved.
 Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature.
@@
 At high pressure, melting temperature substantially depends on pressure (see Ref.
-Trachenko (2024) for the functional form of melting lines).
-At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) set by FPCs in Eq.
-(4) because microscopically, melting involves breaking chemical bonds.
-Experimentally, melting temperatures correlate with cohesive energies and are on the order of $\omega_{\rm D}$ in the range 102-103 K for many systems Ubbelohde (1978); Grimvall and Sjödin (1974).
-As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs in Eq.
+Trachenko (2024) for the functional form of pressure-temperature melting lines).
+At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds.
+This cohesive energy is set by FPCs in Eq.
 (4).
-This then gives the following role of FPCs in setting the states of matter at environmental conditions.
+Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq.
+(9)).
+As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs.
+This gives the following role of FPCs in setting the states of matter.
 FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions.
@@
 Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only.
-As in the case of the second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
-The overlap of our environmental conditions of pressure and temperature with melting and boiling temperatures set by FPCs gives rise to two states important for life: at our environmental conditions, H2O can exist as water and carbon-based substances exist as bonded solidlike structures.
-Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only at environmental conditions and hence no water-based life.
-Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures substances would not exist as cohesive states but as the gas state only.
+We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
+The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life.
+First, H2O can exist as liquid water at our typical planetary conditions.
+Second, carbon-based substances can exist as bonded solidlike structures.
+Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only and hence no water-based life.
+Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures would not exist as cohesive states but as the gas states only.
 We note that classical entropies of solids and gases contain FPCs explicitly.
@@
 $S_{g}\propto\ln\frac{\hbar^{3}m_{p}^{\frac{3}{2}}}{e^{6}m_{e}^{3}}$ (16)
+The same FPCs interestingly feature in both numerators and denominators in Eqs.
+(14) and (16).
 $S_{g}$ in Eq.
-(16) is more sensitive to the variation of FPCs as compared to $S_{s}$ in Eq.
-(14) so that larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster (note that larger $\hbar$ and smaller $m_{e}$ and $e$ decrease the cohesive energy in Eq.
-(4), favouring gases over solids).
-Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective, however this may change if the same solid can exist as a mixture of the stable and many long-lived metastable states (see Section 3.2 and next section for a more detailed discussion).
+(16) is more sensitive to FPCs as compared to $S_{s}$ in Eq.
+(14): larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster.
+The same variation - larger $\hbar$ and smaller $m_{e}$ and $e$ also decrease the cohesive energy in Eq.
+(4), favouring gases over solids.
+Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
 Differently to solid and gas entropies we considered, no closed equation for the liquid entropy exists which corresponds to experimental properties of real liquids such as $c_{v}=3$ close to melting ($c_{v}$ is the constant-volume specific heat).
@@
 3.5 Metastable states and chemical reactions
-In three previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions.
-The usual statistical-mechanical treatment of these transitions and phases assumes equilibrium phases corresponding to the global minimum of the free energy.
-However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the minimum of the free energy.
-Nevertheless, metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition point corresponds to the equality of the free energies of both phases (see Ref.
+In previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions.
+The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy.
+However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy.
+Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref.
 Brazhkin (2006) for review).
 When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
-Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the free energy minimum for a given composition.
+Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition.
 The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006).
-As in the case of quartz-cristobalite mixture discussed in Section 3.2, the ability of a given substance to exist in many long-lived metastable states increases the phase space and entropy.
 Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al.
 .
-Transitions between metastable states involves the same processes as transitions between equilibrium states.
-These processes are governed by the equality of the free energy of long-lived metastable states at the transition point.
-Hence, similarly to what we find for equilibrium transitions in earlier sections, the observability of transitions betweeen different metastable states and hence the existence of numerous metastable phases is set by FPCs.
-Similarly to what we saw in Section 3.2, the phase space and entropy at a given temperature increase due to many long-lived metastable states emerging as a result of these transitions.
+Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs.
+As in the equilibrium case, these transitions correspond to the free energy crossover.
+Hence, similarly to what we saw for equilibrium transitions in earlier sections, the observability and operation of transitions between different metastable states and hence the existence of numerous metastable phases is set by FPCs.
+Similarly to what we saw in Section 3.2, the entropy increases due to new metastable states emerging as a result of these transitions because the states formed at higher temperature tend to be more symmetric with more open structure and have larger phase space as a result (this larger phase space and higher entropy become frozen in the low-temperature metastable phase).
 The same assertion of observability of transitions and different phases applies to chemical reactions and reactions products.
@@
 We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019).
-A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and symmetry Zee (2007).
+A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007).
 FPCs would then be calculable from this theory.
-In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the phase space and entropy.
-Metastable states were important in this discussion because at given environmental parameters (e.g., temperature), the collection of many long-lived metastable states (as opposed to just one thermodynamically stable phase) give the increase of the phase space and entropy.
-We discussed numerous metastable states in condensed matter physics and chemistry in Section 3.
-However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales of matter in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the current age of the Universe Brazhkin (2024)).
-These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading new states such as the nuclear synthesis in stars).
+In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the entropy.
+Long-lived metastable states were important in this discussion because at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy.
+However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)).
+These long-lived metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars.
+In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
 We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and long-lived metastable phases emerging, with the associated increase of the phase space and entropy.
@@
 One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy.
-If most of the current entropy is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
+If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
 $S_{BH}=4\pi\frac{G}{\hbar c}M_{BH}^{2}$ (17)
@@
 However, small $\hbar$ and $c$ give large $\alpha$ in Eq.
-(2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion of smaller BHs).
+(2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
 This discussion and the conjecture of FPCs acting to increase the entropy need further study.
-This includes detailing entropy changes in processes at different levels of structure organisation in the Universe, identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and refining our models setting constraints on FPCs Adams (2019).
+This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on.
 Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
@@
 We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge.
-This implies that the current values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
+This suggests that the values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
 This list of effects and phenomena in condensed matter can be probably extended significantly.
@@
 This and similar insights can provide useful links between different areas of physics and open up new discussions.
-Last by not least, deeper understanding the role of FPCs in condensed matter physics is important on its own.
-Understanding lower and upper bounds of different condensed matter properties including viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al.
+Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own.
+A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models.
+These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al.
 (2017); Sommer et al.
@@
 (2014); Trotzky et al.
-(2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019), thermal conductivity in insulators Trachenko et al.
+(2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al.
 (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al.
-(2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) has deepened our understanding of how these properties operate and benefited our physical models of these properties.
-Understanding how FPCs govern the observability and operation of entirely new effect and phenomena is another useful insight enhancing our understanding of condensed matter physics.
+(2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and superconductivity Trachenko et al.
+.
+Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.
 I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.

### v3

@@
 In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena.
-We discuss first and second-order phase transitions and emerging new states including long-lived metastable states.
-We also discuss transitions between metastable states, chemical reactions and their products.
-These examples show that different values of FPCs would disable these transitions, and new phases as a result of these transitions would not emerge.
-This suggests that the current values of FPCs, by enabling transitions which give rise to new states, act to increase the phase space and entropy.
+We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling.
+We also discuss metastable states, transitions between them, chemical reactions and their products.
+An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs.
+We show that the new states emerging as a result of various transitions increase the phase space and entropy.
+Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge.
+Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
 1 Introduction
-Fundamental physical constants (FPCs), determine physical processes at all energy and length scales of our experiments and theories.
+Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories.
 These different length scales are illustrated in Figure 1.
@@
 Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
-The top in Figure 1 involves cosmological length scales and a balance between the gravitational constant and the cosmological constant.
+The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant.
 This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
@@
 This has deepened our understanding of these properties and their physical models.
-In this Perspective, we discuss a wider role of FPCs in condensed matter: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects.
+In this Perspective, we discuss a wider role of FPCs: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects.
 In the next section, we discuss why discussing FPCs at the condensed matter scale is important.
-This is followed by several examples in condensed matter physics showing how FPCs set observability and operation of phase transitions, the existence of different phases and states of matter emerging as a result of these transitions, including numerous metastable states.
-We also discuss chemical reactions and their products.
-These examples show that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge.
-This suggests that the current values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
+This is followed by considering second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling.
+This discussion shows how FPCs set the observability and operation of phase transitions and hence the existence of different states emerging as a result of these transitions, including metastable states.
+An interesting byproduct of this discussion is that the order of magnitude of the transition temperature in, for example, structural phase transitions, can be calculated from FPCs.
+We show that the new states emerging from various transitions increase the phase space and entropy.
+We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states due to these transitions would not emerge.
+Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
 This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
@@
 In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
-3 Observability and operation of phase and other transitions. Existence of different phases and states
+3 Observability and operation of phase and other transitions. Emergence of new states and entropy
 3.1 Fundamental and environmental parameters
 We start with separating two types of parameters.
-The first is the set of FPCs.
-The second type are the environmental parameters such as, for example, temperature and pressure at typical planetary conditions where we do our observations.
-These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the temperature of the planet.
-In several examples in this section, we consider a certain range of environmental parameters (for example, the temperature range 102-103 K at typical planetary conditions in the Solar system) and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
+The first type are FPCs.
+The second type are the environmental parameters such as temperature and pressure at typical planetary conditions where we do our observations.
+These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature.
+In several examples in this section, we consider a certain range of environmental parameters ($\Delta T$,$\Delta P$) around, for example, temperature (102-103 K) and pressure range at typical planetary conditions in the Solar system and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
 We will be using and discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$.
@@
 where $c$ is the speed of light in vacuum.
-$\alpha$ and $\beta$ are of particular importance when discussing processes at different length scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on.
+$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on.
 These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
-It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter states, solids and liquids, are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
+It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
 $a_{\rm B}=\frac{4\pi\epsilon_{0}\hbar^{2}}{m_{e}e^{2}}$ (3)
@@
 $E_{\rm R}=\frac{m_{e}e^{4}}{32\pi^{2}\epsilon_{0}^{2}\hbar^{2}}$ (4)
-$a_{\rm B}$ is about 0.5 Å and gives a shorter range of typical interatomic separations condensed phases in the range 1-3 Å.
-$E_{\rm R}$ is about 13.6 eV and gives a larger value of cohesive energy in condensed phases in the typical range 1-10 eV.
+$a_{\rm B}\approx 0.5$ and is on the short side of interatomic separations in condensed phases in the typical range 1-3 Å.
+$E_{\rm R}\approx$ 13.6 eV and is on the large side of cohesive energies in condensed phases in the range 1-10 eV.
 3.2 Second-order phase transitions
-Our first case study are the second-order phase structural transitions in solids.
-We consider common silica SiO2 and its low-temperature phase, quartz.
+3.2.1 Enropy increase due to a phase transition
+Our first case study are second-order phase structural transitions in solids.
+Let’s consider common silica SiO2 and its low-temperature phase, quartz.
 Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions.
-Let’s first consider $\beta$-cristobalite, the high-temperature phase emerging at the phase transition around 1800 K Hammonds et al.
+As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition Hammonds et al.
 (1996); Dove (1997).
@@
 $\Delta\Gamma=\frac{1}{(2\pi\hbar)^{3N}}\prod\limits_{i}\Delta p_{i}\Delta q_{i}$ (6)
-Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq.
-(6) and taking the relation between $\Delta\Gamma$ and $S$ as
+Using $\langle\Delta p_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ and $\langle\Delta q_{i}\rangle^{2}\propto\frac{T}{\omega_{i}^{2}}$ for each phonon mode in Eq. (6) and
 $S=\ln\Delta\Gamma$ (7)
@@
 This increases $\Delta\Gamma$ and $S$ of cristobalite compared to quartz.
-Cristobalite is a long-lived metastable phase of silica which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al.
+Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$.
+Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
+$\tilde{\omega}_{i}^{2}=\omega_{i}^{2}+\frac{1}{2}\delta_{i}\eta^{2}$ (8)
+where $\omega_{i}$ are harmonic frequencies and $\delta_{i}$ are the anharmonicity coefficients.
+In the high-temperature phase above the phase transition temperature $T_{c}$, $\eta=0$. Phonon frequencies become smaller as a result, increasing the entropy.
+Importantly, the entropy increase is ultimately related to a phase transition which gives rise to the second high-entropy phase.
+Without a phase transition operating, the entropy would increase according to Eq.
+(5) within the single low-temperature phase only but without larger entropy coming from the second high-temperature phase.
+The entropy increase due to the phase transition can take place both above and below $T_{c}$.
+For reversible phase transition where the low-temperature phase is always recovered below $T_{c}$, the entropy increase takes place above $T_{c}$.
+In our model quartz-cristobalite example, this gives
+$S_{C}(T>T_{c})>S_{Q}(T>T_{c})$ (9)
+where $S_{C}(T>T_{c})$ is the entropy of the high-temperature cristobalite phase above $T_{c}$ and $S_{Q}(T>T_{c})$ is the entropy that quartz would have above $T_{c}$ without the phase transition operating.
+More generally, (9) applies to any two low and high-temperature phases.
+The entropy increase can also take place below $T_{c}$ if the high-temperature phase becomes a long-lived metastable state below $T_{c}$.
+In our model quartz-cristobalite example, the phase transition is a reconstructive phase transition with large energy barriers separating the two phases.
+As a result, cristobalite is a long-lived metastable phase which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al.
 (1996); Dove (1997).
@@
 Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems.
-These functions are defined in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970).
+These functions are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970).
 This is discussed in Section 3.5 in more detail.
-As a more general case compared to pure metastable cristobalite, let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below the transition point where both phases co-exist, $S_{Q-C}$.
+We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$.
 This mixture consists of any proportion of the two phases with the total number of atoms $N$.
@@
 Hence,
-$S_{Q-C}>S_{Q}$ (8)
-We therefore find that the entropy increase (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase.
-The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state at low $T$.
-Without a phase transition operating in the considered range, the entropy would increase according to Eq.
-(5) within the single quartz phase but without the higher entropy contribution from the second metastable phase.
+$S_{Q-C}(T<T_{c})>S_{Q}(T<T_{c})$ (10)
+The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state below $T_{c}$.
+The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
+Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition.
+In a set of solids, phase transitions result in larger phase space and higher entropy.
+Shown minerals are for illustration only.
+3.2.2 Observability of phase transitions. Entropy without a phase transition
 We now observe that if we consider different values of FPCs, the phase transition would not operate.
-The upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants Trachenko et al.
+The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants Trachenko et al.
 .
 This bound follows from combining Eq.
-(4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ as
+(4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ Trachenko and Brazhkin (2020)
 $\frac{\hbar\omega_{m}}{E}=\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (11)
-where $m$ is the atom mass.
-Setting $E=E_{\rm R}$ gives
+where $m$ is the atom mass. Setting $E=E_{\rm R}$ gives
 $\omega_{m}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (12)
-which has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
+$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
 $\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (13)
@@
 (12)-(13).
-Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs.
+Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs.
 (1) and (2) unchanged and hence maintain all preceding essential processes (e.g.
 those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
-Larger $\omega_{m}$ implies larger $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered.
-This, in turn, implies that the high-temperature classical result (5) is no longer applicable because the system is now in the quantum regime.
+Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered.
+This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime.
 In this regime, instead of Eq.
@@
 $S_{s}$ is close to 0 when $T\ll\hbar\omega_{\rm D}$.
-This means that the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition.
+Then the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition.
 The phase transition becomes inoperative.
-Therefore, the current values of fundamental constants, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy in Eq. (8).
-A binary quartz-cristobalite mixture with increased entropy (8) is readily generalised to a collection of many systems existing in long-lived metastable states.
-Figure 2 illustrates this point.
-The increase of the phase space and entropy in these metastable states is generally expected because (a) high-temperature phases tend to be more symmetric Landau and Lifshitz (1970) with more open structure and have a larger number of soft phonons as a result Hammonds et al.
-(1996); Dove (1997) and (b) the increased phase space and entropy in the high-temperature phases become “frozen-in” in metastable states at low $T$.
-Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition.
-The mixture of quartz and metastable cristobalite at low temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms).
-Phase transitions between different states enable sets of long-lived metastable systems with higher combined entropy at the same temperature.
-Shown minerals are for illustration only.
-This case study gives us several insights.
+The observability of phase transitions can also be discussed on the basis of $T_{c}$.
+In phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997).
+This energy is on the order of the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$.
+Using Eq.
+(12), this gives
+$T_{c}\approx\hbar{\omega_{m}}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{2}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (15)
+To estimate $T_{c}$, we write $m=Am_{p}$, where $A$ is the atomic mass. Using Eqs. (12) and (13), $T_{c}$ in Eq. (15) is
+$T_{c}=\frac{\hbar\omega_{m}^{u}}{A^{\frac{1}{2}}}$ (16)
+$A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1 to 15, with an average value of 8.
+Using this value and $\omega_{m}^{u}=3680$ K in Eq.
+(13), Eq.
+(16) gives $T_{c}\approx$ 500 K.
+This is close to $T_{c}$ calculated from typical parameters of the anharmonic phonon model Dove (1993) and is of the right order of magnitude of $T_{c}$ in the range $10^{2}-10^{3}$ K typically observed.
+We see that the order of magnitude of $T_{c}$ (15) is set by FPCs.
+Were different values of FPCs to give $T_{c}$ on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions.
+This is consistent with our earlier argument involving small entropy in Eq.
+(14) due to large $\omega_{\rm D}$.
+From this and previous sections, we conclude that the current values of FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
+3.2.3 Summary
+The discussion in Section 3.2 gives us several insights.
 First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum.
 Second, FPCs affect the observability and operation of entire effects such as phase transitions.
-Third, FPCs, by enabling phase transitions, can increase the phase space and entropy.
+Third, FPCs, by enabling phase transitions, increase the phase space and entropy.
 3.3 Superconducting transition
@@
 Melting and boiling phase transitions would be unobservable.
-Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only.
+Conversely, FPCs resulting in small cohesive energies and hence low melting and boiling temperatures imply that matter would exist in the gas state only.
 We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
@@
 Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
-Differently to solid and gas entropies we considered, no closed equation for the liquid entropy exists which corresponds to experimental properties of real liquids such as $c_{v}=3$ close to melting ($c_{v}$ is the constant-volume specific heat).
-This is due to the fundamental no-small parameter problem of the liquid theory (see Ref.
+Similarly to what we saw in Section 3.2 for the second-order transitions, the first-order phase transitions and the emerging high-temperature gas phase increase the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat).
+Similarly to the second-order transitions, the first-order transitions are enabled by FPCs as discussed earlier in this section.
+Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref.
 Trachenko (2024) for a quick review and Refs.
 Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion).
-Existing models such hard-sphere and Van der Waals models and their extensions give $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically.
-However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: energy of the system is the sum of energies of its excitations Lifshitz and Pitaevskii (2006).
+Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023c).
+However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations Lifshitz and Pitaevskii (2006).
 In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024).
@@
 3.5 Metastable states and chemical reactions
-In previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions.
+In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions.
 The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy.
@@
 Brazhkin (2006) for review).
-When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
+When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
 Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition.
@@
 FPCs would then be calculable from this theory.
-In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the entropy.
-Long-lived metastable states were important in this discussion because at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy.
-However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)).
-These long-lived metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars.
-In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
-We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and long-lived metastable phases emerging, with the associated increase of the phase space and entropy.
+In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy.
+We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy.
 We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy.
@@
 This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
-The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism (see Ref.
-Trachenko (2023d) for the review of related earlier discussions) where entropy increase is analogous to selection due to environmental pressure in biological evolution.
-The mechanism of variability (fluctuations) of FPCs, analogous to genetic mutation in biological evolution, is currently unknown, although several mechanisms were discussed (see, e.g., Refs.
-Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013)).
+We recall that the discussion in Section 3 included long-lived metastable states.
+We saw that at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy.
+As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)).
+These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars.
+In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
+The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution.
+The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g.
+Refs.
+Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013) and Ref.
+Trachenko (2023d) for review).
 Cyclic or oscillatory models of the Universe (see, e.g., Refs.
@@
 Trachenko (2017) at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
-We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space or close to it.
+We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space and entropy or close to it.
 Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g.
@@
 Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles.
-There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (this selection applies to FPCs and FPC-enabled structures existing at the time).
+There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (similarly to biological evolution, this selection applies to FPCs and FPC-enabled structures existing at the time).
 This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence.
@@
 This invites an input from the optimisation theory.
-One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy.
+An example related to this discussion and involving constraints is the black hole (BH) entropy.
 If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
 However, small $\hbar$ and $c$ give large $\alpha$ in Eq.
-(2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
+(2), breaking many constraints required for key processes of structure formation and disrupting these processes Barrow (2003); Adams (2019).
+This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
+We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
 This discussion and the conjecture of FPCs acting to increase the entropy need further study.
-This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on.
-Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
+This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on.
+Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
 5 Discussion and summary

### v4

@@
 Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge.
-Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
+This suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, promote entropy increase.
+We propose that this is a general effect that applies to different levels of structure in the Universe.
 1 Introduction
@@
 The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs.
-Uzan (2003, 2011); Tiesinga et al.
-(2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
+[1, 2, 3, 4]), and their recommended values are updated and maintained in the National Institute for Standards and Technology database [5].
 Figure 1: Different scales of structure in the physical world including the range of typical sizes studied in condensed matter physics.
@@
 In section 3, we discuss a wider role of fundamental constants affecting observability and operation of entire phenomena and effects, including second-order structural and superconducting transitions, first-order melting transition, metastable transitions and chemical and other reactions.
-Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al.
-(2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003).
-The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
-Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006).
-The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983).
+Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future [6, 7, 8, 9, 10, 11, 12, 13, 1].
+The “barcodes of ultimate reality” [6], FPCs also set our Universe apart from those worlds where these constants might be different.
+Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14].
+The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental [15].
 We do not know what kind of theoretical tools we need to address the problem.
 It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
-A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al.
-(2022); Cahn (1996); Hogan (2000); Adams (2019)?
+A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed [6, 7, 8, 9, 10, 11, 12, 13]?
 This is done on the basis of our physical models.
@@
 According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms.
-Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
+Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks [16, 12].
 The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant.
-This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
+This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
 For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe.
-This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019).
+This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13].
 The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are.
 Ref.
-Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC.
+[13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC.
 It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only).
@@
 It was realised that FPCs set upper or lower bounds on several key condensed matter properties.
-Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al.
-(2017); Sommer et al.
-(2011); Bardon et al.
-(2014); Trotzky et al.
-(2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al.
-(2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al.
-(2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b) and superconducting temperature Trachenko et al.
-.
+Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30], velocity gradients that can be set up in cells using biochemical energy [31] and superconducting temperature [32].
 This has deepened our understanding of these properties and their physical models.
@@
 We show that the new states emerging from various transitions increase the phase space and entropy.
-We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states due to these transitions would not emerge.
-Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
-This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
+We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new statess would not emerge.
+This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy.
+We propose that this is a general effect that applies to other levels of structure in the Universe.
+This discussion invites an analogy with an evolutionary process where entropy increase selects for FPCs.
 2 Why filling the condensed matter gap is important
@@
 First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
-Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs Adams (2019).
+Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs [13].
 On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories.
@@
 The first type are FPCs.
-The second type are the environmental parameters such as temperature and pressure at typical planetary conditions where we do our observations.
-These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature.
-In several examples in this section, we consider a certain range of environmental parameters ($\Delta T$,$\Delta P$) around, for example, temperature (102-103 K) and pressure range at typical planetary conditions in the Solar system and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
-We will be using and discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$.
+The second type are the environmental parameters such as temperature $T$ and pressure $P$ at typical planetary conditions where we do our observations.
+These parameters are not fundamental but are due to historical accidents [33] such as the distance between the planet and its star which affects the planet’s temperature.
+In the Solar system, for example, $T$ is in the range 102-103 K.
+In this section, we will consider how FPCs affect the observability and operation of condensed matter phenomena such as phase transitions and other reactions at our environmental conditions.
+We will be discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$.
 We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
@@
 $\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on.
-These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
+These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong coupling constant) [6, 13].
 It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
@@
 $E_{\rm R}=\frac{m_{e}e^{4}}{32\pi^{2}\epsilon_{0}^{2}\hbar^{2}}$ (4)
-$a_{\rm B}\approx 0.5$ and is on the short side of interatomic separations in condensed phases in the typical range 1-3 Å.
-$E_{\rm R}\approx$ 13.6 eV and is on the large side of cohesive energies in condensed phases in the range 1-10 eV.
+$a_{\rm B}\approx 0.5$ and is on the shorter side of interatomic separations in condensed phases in the typical range 1-3 Å.
+$E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1-10 eV.
 3.2 Second-order phase transitions
@@
 Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions.
-As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition Hammonds et al.
-(1996); Dove (1997).
-The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
+As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition [34, 35].
+The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is [36]:
 $S_{s}(T\gg\hbar\omega_{\rm D})=3N\left(\ln\frac{T}{\hbar\bar{\omega}}+1\right)$ (5)
@@
 Here and below, $k_{\rm B}=1$.
-As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz.
-As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al.
-(1996); Dove (1997).
+As expected for high-temperature phases on general grounds [36], cristobalite has a more symmetric and open structure compared to quartz.
+As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra [34, 35].
 This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq.
@@
 Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$.
-Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
+Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are [35, 37]:
 $\tilde{\omega}_{i}^{2}=\omega_{i}^{2}+\frac{1}{2}\delta_{i}\eta^{2}$ (8)
@@
 More generally, (9) applies to any two low and high-temperature phases.
-The entropy increase can also take place below $T_{c}$ if the high-temperature phase becomes a long-lived metastable state below $T_{c}$.
-In our model quartz-cristobalite example, the phase transition is a reconstructive phase transition with large energy barriers separating the two phases.
-As a result, cristobalite is a long-lived metastable phase which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al.
-(1996); Dove (1997).
-According to Eq.
-(5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity).
-Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems.
-These functions are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970).
+The entropy increase can also take place below $T_{c}$ if the high-temperature phase is a long-lived metastable state below $T_{c}$.
+In our quartz-cristobalite example, the phase transition is a reconstructive transition with large energy barriers separating the two phases.
+As a result, cristobalite is a long-lived metastable phase which can be cooled back to room or any other temperature below $T_{c}$ [34, 35].
+We note here that the entropy and other thermodynamic functions of non-equilibrium metastable systems are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics [36].
 This is discussed in Section 3.5 in more detail.
-We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$.
+We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at any temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$.
 This mixture consists of any proportion of the two phases with the total number of atoms $N$.
 We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$.
-$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier.
+$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of cristobalite is larger as discussed earlier.
 Hence,
 $S_{Q-C}(T<T_{c})>S_{Q}(T<T_{c})$ (10)
-The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state below $T_{c}$.
+The increased phase space and entropy above $T_{c}$ become “frozen-in” in the metastable state below $T_{c}$.
 The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
@@
 We now observe that if we consider different values of FPCs, the phase transition would not operate.
-The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants Trachenko et al.
-.
+The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants [32].
 This bound follows from combining Eq.
-(4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ Trachenko and Brazhkin (2020)
+(4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ [17]
 $\frac{\hbar\omega_{m}}{E}=\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (11)
@@
 $\omega_{m}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (12)
-$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
+$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as [32]
 $\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (13)
 $\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1).
-As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
-Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs.
+As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 [6, 13].
+Let’s consider what would happen to a phase transition were FPCs to take different values.
+Let’s consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs.
 (12)-(13).
@@
 (1) and (2) unchanged and hence maintain all preceding essential processes (e.g.
-those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
+stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
 Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered.
@@
 In this regime, instead of Eq.
-(5), the entropy is Landau and Lifshitz (1970)
+(5), the entropy is [36]
 $S_{s}(T\ll\hbar\omega_{\rm D})=\frac{4\pi^{4}}{5}N\left(\frac{T}{\hbar\omega_{\rm D}}\right)^{3}$ (14)
@@
 The observability of phase transitions can also be discussed on the basis of $T_{c}$.
-In phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997).
-This energy is on the order of the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$.
+Phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), expectedly predict that the order of magnitude of $T_{c}$ is set by the energy of interaction between particles [35].
+This energy is close to the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$.
 Using Eq.
-(12), this gives
+(12),
 $T_{c}\approx\hbar{\omega_{m}}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{2}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (15)
@@
 $T_{c}=\frac{\hbar\omega_{m}^{u}}{A^{\frac{1}{2}}}$ (16)
-$A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1 to 15, with an average value of 8.
-Using this value and $\omega_{m}^{u}=3680$ K in Eq.
-(13), Eq.
+$T_{c}$ can be estimated by recalling that $A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1-15, with an average of 8.
+Using this value and $\omega_{m}^{u}=3680$ K (13), Eq.
 (16) gives $T_{c}\approx 500$ K.
-This is close to $T_{c}$ calculated from typical parameters of the anharmonic phonon model Dove (1993) and is of the right order of magnitude of $T_{c}$ in the range $10^{2}-10^{3}$ K typically observed.
+This is close to $T_{c}$ derived using typical parameters of the anharmonic phonon model and Landau theory [38] and is of the right order of magnitude of $T_{c}$ typically observed in the range $10^{2}$-$10^{3}$ K.
 We see that the order of magnitude of $T_{c}$ (15) is set by FPCs.
-Were different values of FPCs to give $T_{c}$ on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions.
-This is consistent with our earlier argument involving small entropy in Eq.
-(14) due to large $\omega_{\rm D}$.
+Were different values of FPCs to give $\omega_{m}$ and $T_{c}$ in Eq.
+(15) on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions.
+This is consistent with our earlier result: large $\omega_{\rm D}$ gives small entropy in Eq.
+(14) and no phase transition.
 From this and previous sections, we conclude that the current values of FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
@@
 Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity.
-We have recently asked Trachenko et al.
-whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get?
+We have recently asked [32] whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get?
 Fundamental limit here specifically means the limit in terms of FPCs.
 Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq.
-(13) Trachenko et al.
-.
+(13) [32].
 $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$.
@@
 Using Eq.
-(13), this gives $T_{c}^{u}$ as Trachenko et al.
+(13), this gives $T_{c}^{u}$ as [32]
 $T_{c}^{u}=C\omega_{m}^{u}=C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\approx 10^{2}-10^{3}\,\rm{K}$ (17)
@@
 Consistent with Eq.
-(17) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al.
-(2020); Duan et al.
-(2017); Pickard et al.
-(2019).
+(17) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides [39, 40, 41].
 This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above.
@@
 At high pressure, melting temperature substantially depends on pressure (see Ref.
-Trachenko (2024) for the functional form of pressure-temperature melting lines).
-At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds.
+[42] for the functional form of pressure-temperature melting lines).
+At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy [43] because microscopically, melting involves breaking chemical bonds.
 This cohesive energy is set by FPCs in Eq.
 (4).
-Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq.
+Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems [43, 44] (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq.
 (11)).
@@
 First, H2O can exist as liquid water at our typical planetary conditions.
-Second, carbon-based substances can exist as bonded solidlike structures.
-Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only and hence no water-based life.
-Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures would not exist as cohesive states but as the gas states only.
+Second, carbon-based molecules and compounds can exist as cohesive (bonded) structures.
+Much larger cohesive energy due to different values of FPCs would result in H2O existing as a solid only and hence no water-based life.
+Much smaller values of cohesive energies due to different FPCs would result in H2O and carbon-based structures not existing as cohesive states but as gas states only.
 We note that classical entropies of solids and gases contain FPCs explicitly.
-Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq.
+Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ [36], $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq.
 (12) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
 $S_{s}\propto\ln\frac{\hbar^{2}m_{p}^{\frac{1}{2}}}{e^{4}m_{e}^{\frac{3}{2}}}$ (18)
-The entropy of the classical gas is Landau and Lifshitz (1970)
+The entropy of the classical gas is [36]
 $S_{g}=\frac{3}{2}N\left(\ln T+1\right)+N\left(\ln v+\ln\left(\frac{m}{2\pi\hbar^{2}}\right)^{\frac{3}{2}}+1\right)$ (19)
 where $v$ is the volume per atom and $m$ is the atom mass.
-With $v\propto a_{\rm B}^{3}$ and $m\propto m_{p}$,
+Using $v\propto a_{\rm B}^{3}$, Eq. (3) and $m\propto m_{p}$ gives
 $S_{g}\propto\ln\frac{\hbar^{3}m_{p}^{\frac{3}{2}}}{e^{6}m_{e}^{3}}$ (20)
@@
 Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref.
-Trachenko (2024) for a quick review and Refs.
-Trachenko and Brazhkin (2016); Trachenko (2023c) for a more detailed discussion).
-Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023c).
-However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations Lifshitz and Pitaevskii (2006).
-In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023c); Proctor (2020, 2021); Chen (2022); de With (2024).
+[42] for a quick review and Refs.
+[45, 46] for a more detailed discussion).
+Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [45, 46].
+However, the energy, the primary property in statistical physics [36], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations [47].
+In liquids, these excitations are gapped phonons and particle jumps [45, 46, 48, 49, 50, 51].
 The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy.
@@
 Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref.
-Brazhkin (2006) for review).
+[52] for review).
 When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
 Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition.
-The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006).
-Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al.
-.
+The world of organic compounds relevant for life processes is nearly all metastable [52].
+Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) [53].
 Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs.
@@
 4 Fundamental constants and entropy increase
-We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019).
-A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007).
+We consider that understanding the current values of FPCs requires the discussion of their variation.
+Theories and observations of this variation were previously discussed [54, 55, 56, 57, 58, 59, 60, 6, 13, 61].
+Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13].
+A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) [33] and guided by a unique design and an exquisite symmetry [62].
 FPCs would then be calculable from this theory.
-In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy.
+In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases at given environmental parameters, increase the entropy.
 We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy.
 We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy.
-This suggests that the current values of fundamental constants act to increase the entropy.
-This insight is general and can be applied to different levels of structure in the Universe in Figure 1.
-This includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed Landau and Lifshitz (1970).
+This suggests that the current values of FPCs give rise to entropy increase.
+We discussed the entropy increase due to FPCs in condensed matter systems, however this increase is a general effect and can be applied to different length and energy scales in the Universe on general grounds.
+This is supported by a more specifc observation.
+The temperature of transitions and reactions at different energy scales is related to the energy of interaction between constituents and the relevant cohesive energy at each scale.
+These energies depend on FPCs, implying that FPCs enable transitions and reactions at those scales as they do in condensed matter.
+Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy.
+Hence we expect that FPCs act to increase the entropy at different levels of structure in the Universe in Figure 1.
+The assertion of the entropy increase includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed [36].
+We also note that the environmental parameters can be variable in some processes.
+For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas.
+This variability can affect transitions and reactions and the properties of emergent structures.
+In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions.
+Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
 We recall that the discussion in Section 3 included long-lived metastable states.
-We saw that at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy.
-As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)).
-These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars.
-In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
+We saw that at given environmental parameters, the collection of these states increase the phase space and entropy.
+As recently reviewed [63], metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe [63]).
+These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which are sensitive to the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
 The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution.
@@
 Refs.
-Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013) and Ref.
-Trachenko (2023d) for review).
+[64, 65, 66, 67, 68] and Ref.
+[61] for review).
 Cyclic or oscillatory models of the Universe (see, e.g., Refs.
-Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref.
-Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle.
+[69, 64, 65, 66, 67, 68, 70] and Ref.
+[71] for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle.
 In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates.
 The size of the oscillating Universe in Ref.
-Trachenko (2017) at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
-We might ask whether the current values of FPCs give structures in the Universe that correspond to the maximised volume of the phase space and entropy or close to it.
-Adams concludes Adams (2019) that the currently observed Universe is not the most abundant and complex possible: somewhat different values of FPCs would result in more complex and abundant structures (e.g.
-more baryons and stars, more efficient nuclear synthesis in stars and so on).
-The evolutionary mechanism (with the entropy increase as a selection principle) would imply that structures observed in each cycle are related to the magnitude of fluctuations of FPCs.
-Large fluctuations would destroy the intricate network of constraints on the values of FPCs needed for different structures in the Universe to exist Adams (2019), resulting in the scarcity of structures and smaller phase space.
-Smaller fluctuations would tend to maintain the degree of continuity of structures in different cycles.
-There is therefore an interplay between the magnitude of fluctuations and selection for the entropy increase within the above constraints (similarly to biological evolution, this selection applies to FPCs and FPC-enabled structures existing at the time).
-This interplay may result in different structures emerging along a sequence of cycles and the entropy of these structures fluctuating along this sequence.
-This process involves maximising a property (entropy) as a function of FPCs but within many constraints on their values required to keep different structures and their phase space intact.
+[70] at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
+We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it.
+Adams concludes [13] that our Universe is not the most abundant and complex possible: somewhat different FPCs would give more complex and abundant structures (e.g.
+more baryons and stars, more efficient nuclear synthesis and so on).
+Another interesting observation in Ref.
+[13] is that our Universe does not lie at the centre of allowed range of FPCs.
+The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations.
+We first note that this selection applies to FPCs and FPC-enabled structures existing at the time (as is the case in biological evolution) rather than to currently observed structures.
+Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space.
+Smaller fluctuations maintain the continuity of structures in different cycles.
+As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along this sequence.
+Third, maximising a property (entropy) as a function of FPCs is subject to many constraints on their values.
 This invites an input from the optimisation theory.
 An example related to this discussion and involving constraints is the black hole (BH) entropy.
-If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
+If most of the current entropy in the Universe is associated with the entropy of large BHs [72], one might consider small $\hbar$ and $c$ to maximise the BH entropy
 $S_{BH}=4\pi\frac{G}{\hbar c}M_{BH}^{2}$ (21)
@@
 However, small $\hbar$ and $c$ give large $\alpha$ in Eq.
-(2), breaking many constraints required for key processes of structure formation and disrupting these processes Barrow (2003); Adams (2019).
-This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
+(2), breaking many constraints required for key processes of structure formation and disrupting these processes [6, 13].
+This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
 We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
 This discussion and the conjecture of FPCs acting to increase the entropy need further study.
-This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on.
-Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
+This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [36] and so on.
+Given the scale of the problem we face in understanding the origin of fundamental constants [6, 15, 14], making new observations and asking new questions is useful for progress.
 5 Discussion and summary
 Approaching the challenge to explain fundamental constants started with understanding the role these constants play.
-This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al.
-(2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003).
+This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures [6, 7, 8, 9, 10, 11, 12, 13, 1].
 Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties.
@@
 In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref.
-Adams (2019) for a review of these challenges).
+[13] for a review of these challenges).
 If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe.
-The insight that FPCs act to increase the entropy is one such example.
+The insight that FPCs promote entropy increase is one such example.
 This and similar insights can provide useful links between different areas of physics and open up new discussions.
@@
 A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models.
-These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al.
-(2017); Sommer et al.
-(2011); Bardon et al.
-(2014); Trotzky et al.
-(2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al.
-(2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al.
-(2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and superconductivity Trachenko et al.
-.
+These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30] and superconductivity [32].
 Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.

### v5

@@
-Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy
+Fundamental physical constants, operation of physical phenomena and entropy increase
 Abstract
@@
 More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties.
-In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena.
-We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling.
+Here, we discuss a much wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena.
+We discuss structural and superconducting phase transitions and transitions between different states of matter, with implications for life processes.
 We also discuss metastable states, transitions between them, chemical reactions and their products.
-An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs.
+A byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs only.
 We show that the new states emerging as a result of various transitions increase the phase space and entropy.
@@
 This suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, promote entropy increase.
-We propose that this is a general effect that applies to different levels of structure in the Universe.
+Based on this entropy increase and the associated increase of statistical probability, we conjecture that entropy increase is a selection principle for FPCs considered to be variable in earlier discussions.
 1 Introduction
@@
 [1, 2, 3, 4]), and their recommended values are updated and maintained in the National Institute for Standards and Technology database [5].
-Figure 1: Different scales of structure in the physical world including the range of typical sizes studied in condensed matter physics.
-Fairly recently, fundamental physical constants were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli and velocity gradients that can be set up in cells using biochemical energy (see discussion in the Introduction).
-In section 3, we discuss a wider role of fundamental constants affecting observability and operation of entire phenomena and effects, including second-order structural and superconducting transitions, first-order melting transition, metastable transitions and chemical and other reactions.
+Figure 1: Different scales of structure in the Universe including the typical size range in condensed matter physics.
+Fairly recently, FPCs were shown to put lower and upper bounds on different condensed matter properties - properties that are already observed - including viscosity, diffusion and spin dynamics, electron and heat transport, thermal conductivity in insulators and conductors, the speed of sound, elastic moduli, flow in cells, vibration frequency in solids and superconducting temperature (see Introduction).
+Here, we discuss a wider role of FPCs affecting observability and operation of entire phenomena and effects, including structural and superconducting transitions, transitions between different states of matter and chemical reactions.
 Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future [6, 7, 8, 9, 10, 11, 12, 13, 3].
@@
 Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14].
-The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental [15].
+The problem of understanding FPCs is not that it involves very hard technical calculations but that we do not know anything more fundamental [15].
 We do not know what kind of theoretical tools we need to address the problem.
@@
 Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks [16, 12].
-The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant.
+The top in Figure 1 involves cosmological scales and a balance between the gravitational and cosmological constants.
 This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
-For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe.
+For the rest of intermediate length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe.
 This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13].
@@
 [13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC.
-It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only).
+FPCs involved in different processes at intermediate scales are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from FPCs operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only).
 This means that the parameter space of these FPCs can be better constrained and understood.
@@
 It was realised that FPCs set upper or lower bounds on several key condensed matter properties.
-Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30], velocity gradients that can be set up in cells using biochemical energy [31] and superconducting temperature [32].
+Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli including those in low-dimensional systems [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34].
 This has deepened our understanding of these properties and their physical models.
-In this Perspective, we discuss a wider role of FPCs: extending beyond setting bounds on properties that are already observed, fundamental constants affect observability and operation of entire physical phenomena and effects.
+Here, we discuss a much wider role of FPCs in condensed matter physics: extending beyond setting bounds on properties that are already observed, FPCs affect observability and operation of entire physical phenomena and effects.
 In the next section, we discuss why discussing FPCs at the condensed matter scale is important.
-This is followed by considering second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling.
-This discussion shows how FPCs set the observability and operation of phase transitions and hence the existence of different states emerging as a result of these transitions, including metastable states.
-An interesting byproduct of this discussion is that the order of magnitude of the transition temperature in, for example, structural phase transitions, can be calculated from FPCs.
+This is followed by considering phase transitions including structural and superconducting transitions and transitions between different states of matter with implications for life processes.
+This shows how FPCs set the operation of entire physical phenomena such as phase and other transitions and hence the existence of different states of matter emerging as a result of these transitions.
+An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs only.
 We show that the new states emerging from various transitions increase the phase space and entropy.
-We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new statess would not emerge.
+We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states would not emerge.
 This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy.
-We propose that this is a general effect that applies to other levels of structure in the Universe.
-This discussion invites an analogy with an evolutionary process where entropy increase selects for FPCs.
+We propose that this is a general effect operating at different levels of structure in the Universe.
+The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution.
+Based on the entropy increase and the associated increase of statistical probability, we conjecture that the entropy increase is a selection principle for FPCs considered to be variable in earlier discussion [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33].
 2 Why filling the condensed matter gap is important
-There are several reasons why understanding the role of FPCs at condensed matter length and energy scales and filling the condensed matter gap in Figure 1 is important.
+There are several reasons why understanding the role of FPCs at condensed matter physics scales and filling the condensed matter gap in Figure 1 is important.
 First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
-Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs [13].
+Second, the uncertainty and insufficient development of physical models related to higher-energy processes limits our understanding of FPCs [13].
 On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories.
@@
 It also provides links and relationships to other areas of physics.
-The fourth and final reason discussed in this Perspective is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales.
+The fourth and final reason discussed here is that if we better understand the operation of FPCs in condensed matter physics, we may find that the underlying principles are quite general and can be applied to other length and energy scales.
 This can lead to new insights related to the origin of FPCs and their values.
-In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
-3 Observability and operation of phase and other transitions. Emergence of new states and entropy
+In the next sections, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
+3 Observability and operation of phase transitions. Emergence of new states with higher entropy
 3.1 Fundamental and environmental parameters
-We start with separating two types of parameters.
-The first type are FPCs.
-The second type are the environmental parameters such as temperature $T$ and pressure $P$ at typical planetary conditions where we do our observations.
-These parameters are not fundamental but are due to historical accidents [33] such as the distance between the planet and its star which affects the planet’s temperature.
-In the Solar system, for example, $T$ is in the range 102-103 K.
-In this section, we will consider how FPCs affect the observability and operation of condensed matter phenomena such as phase transitions and other reactions at our environmental conditions.
-We will be discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$.
+[…]
+We will be discussing FPCs central to condensed matter physics: the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$.
+As discussed in this section, the first three constants set the characteristic scales of distance and energy in condensed matter.
+$m_{p}$ sets the scale of lattice oscillation frequency (see section 3.2.2).
 We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
@@
 and the fine-structure constant
-$\alpha=\frac{e^{2}}{\hbar c}$ (2)
+$\alpha=\frac{1}{4\pi\epsilon_{0}}\frac{e^{2}}{\hbar c}$ (2)
 where $c$ is the speed of light in vacuum.
@@
 $E_{\rm R}=\frac{m_{e}e^{4}}{32\pi^{2}\epsilon_{0}^{2}\hbar^{2}}$ (4)
-$a_{\rm B}\approx 0.5$ and is on the shorter side of interatomic separations in condensed phases in the typical range 1-3 Å.
-$E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1-10 eV.
-3.2 Second-order phase transitions
-3.2.1 Enropy increase due to a phase transition
-Our first case study are second-order phase structural transitions in solids.
-Let’s consider common silica SiO2 and its low-temperature phase, quartz.
-Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions.
-As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition [34, 35].
+$a_{\rm B}\approx 0.5$ and is on the shorter side of interatomic separations in condensed phases in the typical range 1–3 Å.
+$E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1–10 eV.
+3.2 Structural phase transitions
+3.2.1 Entropy increase due to a phase transition
+Our first case study are structural phase transitions in solids.
+Let’s consider common silica SiO2 which has several polymorphs at different pressure and temperature, separated by phase transitions.
+As a model example, let’s consider two silica phases: quartz and high-temperature $\beta$-cristobalite phase [44, 45].
 The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is [46]:
 $S_{s}(T\gg\hbar\omega_{\rm D})=3N\left(\ln\frac{T}{\hbar\bar{\omega}}+1\right)$ (5)
-where $\bar{\omega}$ is the average phonon frequency defined as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies and $N$ is the number of atoms.
+where $\bar{\omega}$ is the average phonon frequency set as $\bar{\omega}^{3N}=\prod\limits_{i}\omega_{i}$, $\omega_{i}$ are $3N$ phonon frequencies, $N$ is the number of atoms and anharmonic corrections are assumed small and not considered below.
 Here and below, $k_{\rm B}=1$.
@@
 (5) within the single low-temperature phase only but without larger entropy coming from the second high-temperature phase.
-The entropy increase due to the phase transition can take place both above and below $T_{c}$.
+We consider $T_{c}$ to be within the range of environmental temperature, $T_{1}<T_{c}<T_{2}$.
+The entropy increase due to phase transitions can take place both above and below $T_{c}$.
 For reversible phase transition where the low-temperature phase is always recovered below $T_{c}$, the entropy increase takes place above $T_{c}$.
@@
 We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$.
-$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of cristobalite is larger as discussed earlier.
+As discussed earlier, the entropy of cristobalite is higher than that of quartz.
 Hence,
 $S_{Q-C}(T<T_{c})>S_{Q}(T<T_{c})$ (10)
-The increased phase space and entropy above $T_{c}$ become “frozen-in” in the metastable state below $T_{c}$.
+[…]
 The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
@@
 Shown minerals are for illustration only.
-3.2.2 Observability of phase transitions. Entropy without a phase transition
+3.2.2 Observability of phase transitions. Entropy changes.
 We now observe that if we consider different values of FPCs, the phase transition would not operate.
@@
 (12)-(13).
-Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs.
+Note that this question is well-posed because it is possible to vary $\omega_{m}^{u}$ by varying $m_{e}$ and $m_{p}$ or $\hbar$ and $e$ while keeping $\beta$ and $\alpha$ in Eqs.
 (1) and (2) unchanged and hence maintain all preceding essential processes (e.g.
 stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
-Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered.
+[…]
+$T_{2}\ll\hbar\omega_{m}$ (14)
+Recalling that $\omega_{m}$ is close to $\omega_{\rm D}$, condition (14) implies $T\ll\hbar\omega_{\rm D}$.
 This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime.
@@
 Then the second term in the free energy $F=E-TS$ is close to 0 too, and $F$ now does not have a crossover corresponding to a phase transition.
-The phase transition becomes inoperative.
+The phase transition becomes inoperative, and there is no associated entropy increase discussed in the previous section.
 The observability of phase transitions can also be discussed on the basis of $T_{c}$.
@@
 This is consistent with our earlier result: large $\omega_{\rm D}$ gives small entropy in Eq.
-(14) and no phase transition.
-From this and previous sections, we conclude that the current values of FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
-3.2.3 Summary
-The discussion in Section 3.2 gives us several insights.
-First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum.
-Second, FPCs affect the observability and operation of entire effects such as phase transitions.
-Third, FPCs, by enabling phase transitions, increase the phase space and entropy.
+(15), no phase transition and no associated entropy increase.
+We therefore find that FPCs, by enabling the operation of phase transitions which give rise to new states, increase the phase space and entropy.
+We note that phase transitions in the phase-changing subsystem (PCS) involve the heat flow from another, heat-supplying subsystem (HSS) in order to reach $T_{c}$ and the transition (if the second high-entropy phase in the PCS is metastable below $T_{c}$, the heat can be reversibly returned from the PCS to HSS below $T_{c}$, retaining the entropy increase in the metastable phase and involving little or no net heat exchange between PCS and HSS).
+The PCS and HSS are open, however the combined PCS and HSS subsystem can be considered closed.
+The entropy increase due to phase transitions in the PCS applies to this closed subsystem and represents the entropy increase which is extra as compared to the case where no new phases emerge in the PCS and the entropy increase is due to the trivial heat exchange between the PCS and HSS only.
+In summary, the discussion in Section 3.2 has so far given us several insights:
+  1. 1.
+    FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum.
+  2. 2.
+    FPCs set the observability and operation of entire physical phenomena such as phase transitions (more examples are in Sections 3.3, 3.4 and 3.5).
+  3. 3.
+    FPCs, by enabling phase transitions, increase the phase space and entropy (more discussion is in Sections 3.4, 3.5 and 4 below).
 3.3 Superconducting transition
@@
 (18), we ask how would the observability of superconductivity change were fundamental constants to take different values?
-Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved.
+Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved at our experimental conditions.
 Another interesting realisation comes from observing that a variation of FPCs giving $T_{c}^{u}$ of, for example, 105 K or higher, would imply that many materials would have $T_{c}$ above room temperature.
@@
 We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
-3.4 First-order phase transitions: solids, liquids, gases
+3.4 Transitions between states of matter: solids, liquids and gases
 Three basic states of matter, solids, liquids and gases, are represented by a phase diagram with three first-order phase transition lines as shown in Figure 3.
 In this section, we discuss how the observability of different states of matter and transitions between those states at our environmental conditions is related to FPCs.
-Figure 3: Three states of matter, solids, liquids and gases, separated by three phase transition lines. The triple and critical points are shown.
+Figure 3: Three states of matter (solids, liquids and gases) separated by three phase transition lines. Triple and critical points are shown.
 The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase.
@@
 Conversely, FPCs resulting in small cohesive energies and hence low melting and boiling temperatures imply that matter would exist in the gas state only.
-We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
-The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life.
-First, H2O can exist as liquid water at our typical planetary conditions.
-Second, carbon-based molecules and compounds can exist as cohesive (bonded) structures.
+We see that, similarly to structural phase transitions discussed in Section 3.2, the current values of FPCs set the observability of transitions between different states of matter in Figure 3 and the existence of these states.
+The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life:
+  1. 1.
+    H2O can exist as liquid water at our typical planetary conditions.
+  2. 2.
+    Carbon-based molecules and compounds can exist as cohesive (bonded) structures.
 Much larger cohesive energy due to different values of FPCs would result in H2O existing as a solid only and hence no water-based life.
 Much smaller values of cohesive energies due to different FPCs would result in H2O and carbon-based structures not existing as cohesive states but as gas states only.
+A more detailed consideration shows an additional and more specific role of FPCs in life processes: FPCs have a biofriendly window constrained by biofriendly viscosity and diffusion setting the motion in essential processes in and across cells [32, 33].
 We note that classical entropies of solids and gases contain FPCs explicitly.
@@
 Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
-Similarly to what we saw in Section 3.2 for the second-order transitions, the first-order phase transitions and the emerging high-temperature gas phase increase the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat).
-Similarly to the second-order transitions, the first-order transitions are enabled by FPCs as discussed earlier in this section.
-Differently from solid and gas entropies we considered, discussing liquid entropy is more challenging theoretically due to the fundamental no-small parameter problem of the liquid theory (see Ref.
+Similarly to what we saw in Section 3.2 for structural phase transitions, the emerging high-temperature gas phase increases the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat).
+Similarly to structural transitions, transitions between different states of matter are enabled by FPCs as discussed earlier in this section.
+Differently from solid and gas entropies in Eqs.
+(5), (15) and (20), no equation for the liquid entropy exists due to the fundamental no-small parameter problem of the liquid theory (see Ref.
 [52] for a quick review and Refs.
@@
 Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [55, 56].
-However, the energy, the primary property in statistical physics [36], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy is given by the energies of its excitations [47].
+However, the energy, the primary property in statistical physics [46], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy $E$ is set by the energies of its excitations [57].
 In liquids, these excitations are gapped phonons and particle jumps [55, 56, 58, 59, 60, 61].
-The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}$ is calculated from the energy.
-Nevertheless, liquids are a condensed state of matter, and our discussion of how the interplay of cohesive energies set by FPCs and environmental parameters gives rise to different states of matter applies to liquids too.
+The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}=\frac{dE}{dT}$ (in almost entire range of liquid relaxation time, $E$ and $c_{v}$ are vibrational to a very good approximation [62]).
+Nevertheless, liquids are a cohesive state, and our discussion of how different states of matter emerge depending on cohesive energies set by FPCs applies to liquids too.
 3.5 Metastable states and chemical reactions
-In the previous three sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions.
+In the previous sections, we discussed phase transitions and the existence of different states as a result of these transitions.
 The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy.
-However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy.
-Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref.
+However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest free energy minimum.
+Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref.
 [63] for review).
-When discussing phase transitions in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
+When discussing transitions in these systems, the tacit assumption is the one we mentioned in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
 Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition.
@@
 Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs.
-As in the equilibrium case, these transitions correspond to the free energy crossover.
-Hence, similarly to what we saw for equilibrium transitions in earlier sections, the observability and operation of transitions between different metastable states and hence the existence of numerous metastable phases is set by FPCs.
-Similarly to what we saw in Section 3.2, the entropy increases due to new metastable states emerging as a result of these transitions because the states formed at higher temperature tend to be more symmetric with more open structure and have larger phase space as a result (this larger phase space and higher entropy become frozen in the low-temperature metastable phase).
-The same assertion of observability of transitions and different phases applies to chemical reactions and reactions products.
-Chemical reactions involve changes of chemical bonding whose characteristic energies are set by FPCs.
-Hence these reactions are governed by the interplay of characteristic cohesive energies set by FPCs and environmental parameters.
-For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be unobservable at given environmental parameters.
+Hence, similarly to equilibrium transitions, the observability and operation of transitions between different metastable states and the existence of numerous metastable phases is set by FPCs.
+The entropy of new high-temperature metastable states emerging in these transitions is expected to be higher because these states tend to be more symmetric with more open structure, as is the case for equilibrium transitons.
+The same assertion of observability of transitions and different phases applies to chemical reactions and their products.
+Chemical reactions involve changes of chemical bonds with characteristic energies set by FPCs.
+Hence these reactions are governed by the cohesive energies set by FPCs and environmental parameters.
+For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be inoperable at given environmental parameters.
 4 Fundamental constants and entropy increase
 We consider that understanding the current values of FPCs requires the discussion of their variation.
-Theories and observations of this variation were previously discussed [54, 55, 56, 57, 58, 59, 60, 6, 13, 61].
+Theories and observations of this variation were previously discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33].
 Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13].
@@
 FPCs would then be calculable from this theory.
-In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases at given environmental parameters, increase the entropy.
+In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy.
 We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy.
@@
 This suggests that the current values of FPCs give rise to entropy increase.
-We discussed the entropy increase due to FPCs in condensed matter systems, however this increase is a general effect and can be applied to different length and energy scales in the Universe on general grounds.
-This is supported by a more specifc observation.
-The temperature of transitions and reactions at different energy scales is related to the energy of interaction between constituents and the relevant cohesive energy at each scale.
-These energies depend on FPCs, implying that FPCs enable transitions and reactions at those scales as they do in condensed matter.
+We discussed the entropy increase in condensed matter systems, however this increase applies to different length and energy scales in the Universe on general grounds.
+Recall Eq.
+(16) giving the temperature of phase transitions as a function of FPCs in condensed matter.
+However, the temperature of transitions and reactions depend on FPCs at other energy and length scales as well [6, 13], enabling transitions and reactions at those scales.
 Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy.
-Hence we expect that FPCs act to increase the entropy at different levels of structure in the Universe in Figure 1.
-The assertion of the entropy increase includes the usual proviso that we consider a subsystem or a collection of subsystems in the Universe which are closed or conditionally closed [36].
-We also note that the environmental parameters can be variable in some processes.
+Hence we expect that FPCs promote entropy increase at different levels of structure in the Universe.
+This entropy increase is specifically due to transitions and reactions and comes in addition to the conventional general mechanism of entropy increase (including in the second law of thermodynamics).
+In the conventional mechanism, $S$ of a closed system at a given energy $E_{0}$ is maximised at equilibrium because (a) the statistical probability $d\rho$ is
+$d\rho={\rm const}\cdot\delta(E-E_{0})e^{S}\prod\limits_{i}dE_{i}$ (22)
+where $E_{i}$ are energies of subsystems and (b) $d\rho$ is maximised when $E_{i}$ of all subsystems in their own equilibrium states are equal to their most probable equilibrium values $\bar{E_{i}}$, corresponding to full equilibrium [46].
+As discussed in Section 3.2, the subsystem where transitions and reactions take place (phase-changing subsystem, PCS) is not closed and not in equilibrium because these transitions are enabled by the heat flow from another heat-supplying subsystem.
+However, subsystems can be identified which are (a) larger, include both subsystems and are closed and (b) in equilibrium at long post-reaction times [46].
+The conventional mechanism of increase of $S$ and $d\rho$ applies to these closed subsystems, and our entropy increase in the PCS enabled by FPCs additionally increases the statistical probability in Eq.
+(22).
+As discussed below, identifying collections of subsystems which can be considered as conditionally closed at different levels of the Universe can be of interest in future work.
+We note that the environmental parameters can be variable in some processes.
 For example, nuclear synthesis reactions in a forming star depend on (environmental) pressure and temperature, which, in turn, depend on the mass of collapsing interstellar gas.
 This variability can affect transitions and reactions and the properties of emergent structures.
-In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions.
+In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions themselves.
 Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
-We recall that the discussion in Section 3 included long-lived metastable states.
-We saw that at given environmental parameters, the collection of these states increase the phase space and entropy.
-As recently reviewed [63], metastable states are ubiquitous at different scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe [63]).
-These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which are sensitive to the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
-The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where entropy increase is analogous to selection due to environmental pressure in biological evolution.
-The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, and its operation are unknown, although several proposals were discussed (see, e.g.
-Refs.
-[64, 65, 66, 67, 68] and Ref.
-[61] for review).
+We recall the discussion of long-lived metastable states with higher entropy in Section 3.
+As recently reviewed [66], metastable states are ubiquitous at different levels of matter organization in the Universe including nuclei and astronomical objects.
+The energy barriers setting the lifetime of these states vary in a wide range 10-7 to 109 eV per particle, with relaxation times often exceeding the age of the Universe [66].
+These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which depend on the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
+The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution due to environmental pressure.
+Based on this entropy increase and the associated increase of statistical probability (22), we conjecture that entropy increase is a selection principle for FPCs considered to vary in earlier discussions (see, e.g., Refs.
+[35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33] and references therein).
+In other words, FPCs tend to give the most probable state.
+The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, is unknown, although several proposals were discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33].
+The above conjecture is nevertheless interesting to consider because, if correct, it is analogous to introducing a Darwinian concept of a selection principle, the notion that brought new understanding even though the molecular DNA-based mutation mechanism was not known at the time and came about a century later only.
 Cyclic or oscillatory models of the Universe (see, e.g., Refs.
@@
 In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates.
-The size of the oscillating Universe in Ref.
-[70] at the inversion point is set by the system energy as a parameter in the model and does not need to be small.
+The size of the oscillating Universe at the inversion point in Ref.
+[72] is set by the system energy as a parameter in the model and does not need to be small.
 We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it.
-Adams concludes [13] that our Universe is not the most abundant and complex possible: somewhat different FPCs would give more complex and abundant structures (e.g.
+Adams concludes [13] that the observable Universe is not the most abundant and complex possible: somewhat different values of FPCs would give more complex and abundant structures (e.g.
 more baryons and stars, more efficient nuclear synthesis and so on).
 Another interesting observation in Ref.
-[13] is that our Universe does not lie at the centre of allowed range of FPCs.
+[13] is that our Universe does not lie at the centre of allowed range of values of FPCs.
 The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations.
@@
 Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space.
-Smaller fluctuations maintain the continuity of structures in different cycles.
-As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along this sequence.
-Third, maximising a property (entropy) as a function of FPCs is subject to many constraints on their values.
+On the other hand, smaller fluctuations maintain the continuity of structures in different cycles.
+As a result, different structures can emerge along a sequence of cycles, with the entropy of these structures fluctuating along the sequence.
+Third, it follows that maximising entropy as a function of FPCs is subject to many constraints on their values.
 This invites an input from the optimisation theory.
@@
 However, small $\hbar$ and $c$ give large $\alpha$ in Eq.
-(2), breaking many constraints required for key processes of structure formation and disrupting these processes [6, 13].
+(2), breaking many constraints required for key processes of structure formation and disrupting those processes [6, 13].
 This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
 We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
-This discussion and the conjecture of FPCs acting to increase the entropy need further study.
+The discussion of FPCs acting to increase the entropy and the conjecture of entropy increase as a selection principle for FPCs need further study.
 This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [46] and so on.
@@
 This has enriched condensed matter physics and deepened our understanding of its key phenomena.
-Here, we discussed a wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena.
-We gave examples of operation of first- and second-order phase transitions and emerging new phases, existence of numerous metastable phases, chemical reactions, their products and so on.
+Here, we discussed a much wider role played by FPCs in condensed matter physics: at given environmental parameters, FPCs define observability and operation of entire physical effects and phenomena.
+Our examples included the operation of different phase transitions and emerging new phases, existence of different states of matter, metastable phases, chemical reactions, their products and so on.
+These results were based on a theory of each physical effect and process rather than on a dimensional analysis as might have been construed.
+Ref.
+[31] (pages 479-480) lists important differences between such a theory-based approach and a dimensional analysis.
 We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge.
 This suggests that the values of FPCs, by enabling transitions and new states emerging as a result of these transitions, act to increase the phase space and entropy.
-This list of effects and phenomena in condensed matter can be probably extended significantly.
+The list of effects and phenomena in condensed matter physics discussed here can be probably extended significantly.
 The reason it is important to study how FPCs affect the operation and observability of different effects and phenomena in condensed matter is that this field is developed at the great level of detail.
@@
 [13] for a review of these challenges).
-If the insights into the role of FPCs are found to be general enough, these insights can be applied to structures at other levels in the Universe.
-The insight that FPCs promote entropy increase is one such example.
-This and similar insights can provide useful links between different areas of physics and open up new discussions.
+If the insights into the role of FPCs in condensed matter physics are found to be general enough, these insights can be applied to structures at other levels in the Universe.
+An example is the insight that FPCs promote entropy increase, or the related conjecture of the selection principle for FPCs.
+These and similar insights can provide useful links between different areas of physics and open up new discussions.
 Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own.
 A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models.
-These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30] and superconductivity [32].
+These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34].
 Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.
-I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.
+I am grateful to F. C. Adams, V. V. Brazhkin, M. T. Dove, M. Hutcheon, S. A. Kauffman, Z.-K. Liu, B. Monserrat, A. J. Phillips and C. J. Pickard for discussions and EPSRC for support.

### v6

@@
 Based on this entropy increase and the associated increase of statistical probability, we conjecture that entropy increase is a selection principle for FPCs considered to be variable in earlier discussions.
-1 Introduction
+I Introduction
 Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories.
@@
 The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs.
-[1, 2, 3, 4]), and their recommended values are updated and maintained in the National Institute for Standards and Technology database [5].
+Tiesinga et al.
+(2021); Ashcroft and Mermin (1976); Uzan (2003, 2011)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
 Figure 1: Different scales of structure in the Universe including the typical size range in condensed matter physics.
@@
 Here, we discuss a wider role of FPCs affecting observability and operation of entire phenomena and effects, including structural and superconducting transitions, transitions between different states of matter and chemical reactions.
-Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future [6, 7, 8, 9, 10, 11, 12, 13, 3].
-The “barcodes of ultimate reality” [6], FPCs also set our Universe apart from those worlds where these constants might be different.
-Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14].
-The problem of understanding FPCs is not that it involves very hard technical calculations but that we do not know anything more fundamental [15].
+Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al.
+(2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003).
+The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
+Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006).
+The problem of understanding FPCs is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983).
 We do not know what kind of theoretical tools we need to address the problem.
 It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
-A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed [6, 7, 8, 9, 10, 11, 12, 13]?
+A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al.
+(2022); Cahn (1996); Hogan (2000); Adams (2019)?
 This is done on the basis of our physical models.
@@
 According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms.
-Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks [16, 12].
+Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
 The top in Figure 1 involves cosmological scales and a balance between the gravitational and cosmological constants.
-This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
+This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
 For the rest of intermediate length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe.
-This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13].
+This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019).
 The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are.
 Ref.
-[13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC.
-FPCs involved in different processes at intermediate scales are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from FPCs operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only).
-This means that the parameter space of these FPCs can be better constrained and understood.
+Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC.
 Compared to the above high-energy processes, the discussion of the role of FPCs at the scale of condensed matter has been scarce - this is the scale spanning 108-109 orders of magnitude in Figure 1 or about 20% of the logarithmic scale between the nuclei and the size of Universe.
@@
 It was realised that FPCs set upper or lower bounds on several key condensed matter properties.
-Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli including those in low-dimensional systems [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34].
+Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al.
+(2017); Sommer et al.
+(2011); Bardon et al.
+(2014); Trotzky et al.
+(2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022); Grozdanov (2021), thermal conductivity in insulators Trachenko et al.
+(2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al.
+(2020), elastic moduli including those in low-dimensional systems Brazhkin (2023); Trachenko (2023a), flow in cells Trachenko (2023b, c), vibration frequency in solids and superconducting temperature Trachenko et al.
+.
 This has deepened our understanding of these properties and their physical models.
@@
 We propose that this is a general effect operating at different levels of structure in the Universe.
-The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution.
-Based on the entropy increase and the associated increase of statistical probability, we conjecture that the entropy increase is a selection principle for FPCs considered to be variable in earlier discussion [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33].
-2 Why filling the condensed matter gap is important
+The entropy increase due to FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution.
+Based on this entropy increase and the associated increase of statistical probability, we conjecture that the entropy increase is a selection principle for FPCs considered to be variable in earlier discussions.
+II Why filling the condensed matter gap is important
 There are several reasons why understanding the role of FPCs at condensed matter physics scales and filling the condensed matter gap in Figure 1 is important.
 First and quite generally, having a non-interrupted and continuous set of length and energy scales in Figure 1 over which we can discuss the role of FPCs gives more information and detail to use.
-Second, the uncertainty and insufficient development of physical models related to higher-energy processes limits our understanding of FPCs [13].
+Second, the uncertainty and insufficient development of physical models related to higher-energy processes limits our understanding of FPCs Adams (2019).
 On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories.
@@
 This can lead to new insights related to the origin of FPCs and their values.
-In the next sections, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
-3 Observability and operation of phase transitions. Emergence of new states with higher entropy
-3.1 Fundamental and environmental parameters
+In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
+III Observability and operation of phase transitions. Emergence of new states with higher entropy
+III.1 Fundamental and environmental parameters
+We start with separating two types of parameters.
+The first type are FPCs.
+The second type are the environmental parameters such as temperature $T$ and pressure $P$ at typical planetary conditions where we do our observations.
+These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature.
+In the Solar system, for example, $T$ is in the range 102–103 K.
 […]
-We will be discussing FPCs central to condensed matter physics: the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$.
-As discussed in this section, the first three constants set the characteristic scales of distance and energy in condensed matter.
-$m_{p}$ sets the scale of lattice oscillation frequency (see section 3.2.2).
-We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
-$\beta=\frac{m_{p}}{m_{e}}$ (1)
-and the fine-structure constant
-$\alpha=\frac{1}{4\pi\epsilon_{0}}\frac{e^{2}}{\hbar c}$ (2)
-where $c$ is the speed of light in vacuum.
-$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on.
-These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong coupling constant) [6, 13].
-It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
+FPCs discussed in this paper will include those central to condensed matter physics: the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$.
+These constants define the characteristic scales of distance and energy in condensed matter by setting the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
 $a_{\rm B}=\frac{4\pi\epsilon_{0}\hbar^{2}}{m_{e}e^{2}}$ (1)
@@
 $E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1–10 eV.
-3.2 Structural phase transitions
-3.2.1 Entropy increase due to a phase transition
+We will also discuss the mass of the proton (the composite particle), $m_{p}$, which sets the scale of lattice oscillation frequency (see section III.2.2).
+$m_{p}$ will mostly enter our discussion as the dimensionless ratio
+$\beta=\frac{m_{p}}{m_{e}}$ (3)
+We note another important dimensionless parameter, the fine-structure constant
+$\alpha=\frac{1}{4\pi\epsilon_{0}}\frac{e^{2}}{\hbar c}$ (4)
+where $c$ is the speed of light in vacuum.
+$\alpha$ and $\beta$ play a key role at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on.
+These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong coupling constant) Barrow (2003); Adams (2019).
+$\alpha$ will play an indirect role in our discussion in Section III.2.2 because we don’t consider relativistic effects here.
+This may change if relativistic effects play a role in other condensed matter processes.
+This amounts to the following set of FPCs discussed in this paper:
+${\rm FPCs}=\{\hbar,e,m_{e},\frac{m_{p}}{m_{e}},\alpha\}$ (5)
+III.2 Structural phase transitions
+III.2.1 Entropy increase due to a phase transition
 Our first case study are structural phase transitions in solids.
 Let’s consider common silica SiO2 which has several polymorphs at different pressure and temperature, separated by phase transitions.
-As a model example, let’s consider two silica phases: quartz and high-temperature $\beta$-cristobalite phase [44, 45].
-The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is [46]:
+As a model example, let’s consider two silica phases: quartz and high-temperature $\beta$-cristobalite phase Hammonds et al.
+(1996); Dove (1997).
+The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
 $S_{s}(T\gg\hbar\omega_{\rm D})=3N\left(\ln\frac{T}{\hbar\bar{\omega}}+1\right)$ (6)
@@
 Here and below, $k_{\rm B}=1$.
-As expected for high-temperature phases on general grounds [46], cristobalite has a more symmetric and open structure compared to quartz.
-As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra [44, 45].
+As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz.
+As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al.
+(1996); Dove (1997).
 This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq.
@@
 Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$.
-Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are [45, 47]:
+Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
 $\tilde{\omega}_{i}^{2}=\omega_{i}^{2}+\frac{1}{2}\delta_{i}\eta^{2}$ (9)
@@
 In our quartz-cristobalite example, the phase transition is a reconstructive transition with large energy barriers separating the two phases.
-As a result, cristobalite is a long-lived metastable phase which can be cooled back to room or any other temperature below $T_{c}$ [44, 45].
-We note here that the entropy and other thermodynamic functions of non-equilibrium metastable systems are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics [46].
-This is discussed in Section 3.5 in more detail.
+As a result, cristobalite is a long-lived metastable phase which can be cooled back to room or any other temperature below $T_{c}$ Hammonds et al.
+(1996); Dove (1997).
+We note here that the entropy and other thermodynamic functions of non-equilibrium metastable systems are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970).
+This is discussed in Section III.5 in more detail.
 We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at any temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$.
@@
 […]
-The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
-Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition.
-In a set of solids, phase transitions result in larger phase space and higher entropy.
-Shown minerals are for illustration only.
-3.2.2 Observability of phase transitions. Entropy changes.
+The quartz-cristobalite example is readily generalised to many high-temperature polymorphic phases of different solids emerging as a result of phase transitions, with overall entropy increased.
+III.2.2 Observability of phase transitions. Entropy changes.
 We now observe that if we consider different values of FPCs, the phase transition would not operate.
-The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants [34].
+The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of FPCs Trachenko et al.
+.
 This bound follows from combining Eq.
-(4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ [17]
+(2) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$.
+The ratio involves a harmonic approximation for the relation between $E$ and $\omega_{m}$ and reads Trachenko and Brazhkin (2020); Trachenko (2023a):
 $\frac{\hbar\omega_{m}}{E}=\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (12)
@@
 $\omega_{m}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ (13)
-$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as [34]
+$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
 $\omega_{m}^{u}=\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\approx 3680\,\rm{K}$ (14)
 $\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (3).
-As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 [6, 13].
+As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
 Let’s consider what would happen to a phase transition were FPCs to take different values.
@@
 (3) and (4) unchanged and hence maintain all preceding essential processes (e.g.
-stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
+stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
 […]
@@
 In this regime, instead of Eq.
-(5), the entropy is [46]
+(6), the entropy is Landau and Lifshitz (1970)
 $S_{s}(T\ll\hbar\omega_{\rm D})=\frac{4\pi^{4}}{5}N\left(\frac{T}{\hbar\omega_{\rm D}}\right)^{3}$ (16)
@@
 The observability of phase transitions can also be discussed on the basis of $T_{c}$.
-Phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), expectedly predict that the order of magnitude of $T_{c}$ is set by the energy of interaction between particles [45].
+Phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), expectedly predict that the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997).
 This energy is close to the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$.
@@
 (18) gives $T_{c}\approx 500$ K.
-This is close to $T_{c}$ derived using typical parameters of the anharmonic phonon model and Landau theory [48] and is of the right order of magnitude of $T_{c}$ typically observed in the range $10^{2}$-$10^{3}$ K.
+This is close to $T_{c}$ derived using typical parameters of the anharmonic phonon model and Landau theory Dove (1993) and is of the right order of magnitude of $T_{c}$ typically observed in the range $10^{2}$-$10^{3}$ K.
 We see that the order of magnitude of $T_{c}$ (17) is set by FPCs.
@@
 The entropy increase due to phase transitions in the PCS applies to this closed subsystem and represents the entropy increase which is extra as compared to the case where no new phases emerge in the PCS and the entropy increase is due to the trivial heat exchange between the PCS and HSS only.
-In summary, the discussion in Section 3.2 has so far given us several insights:
+In summary, the discussion in Section III.2 has so far given us several insights:
   1. 1.
@@
   2. 2.
-    FPCs set the observability and operation of entire physical phenomena such as phase transitions (more examples are in Sections 3.3, 3.4 and 3.5).
+    FPCs set the observability and operation of entire physical phenomena such as phase transitions (more examples are in Sections III.3, III.4 and III.5).
   3. 3.
-    FPCs, by enabling phase transitions, increase the phase space and entropy (more discussion is in Sections 3.4, 3.5 and 4 below).
-3.3 Superconducting transition
+    FPCs, by enabling phase transitions, increase the phase space and entropy (more discussion is in Sections III.4, III.5 and IV below).
+III.3 Superconducting transition
 Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity.
-We have recently asked [34] whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get?
+We have recently asked Trachenko et al.
+whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get?
 Fundamental limit here specifically means the limit in terms of FPCs.
 Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq.
-(13) [34].
+(14) Trachenko et al.
+.
 $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$.
@@
 Using Eq.
-(13), this gives $T_{c}^{u}$ as [34]
+(14), this gives $T_{c}^{u}$ as Trachenko et al.
 $T_{c}^{u}=C\omega_{m}^{u}=C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{1}{32\pi^{2}\epsilon_{0}^{2}}\frac{m_{e}e^{4}}{\hbar^{3}}\approx 10^{2}-10^{3}\,\rm{K}$ (19)
@@
 Consistent with Eq.
-(18) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides [49, 50, 51].
+(19) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to current FPCs, we observe superconductivity in the range $T\lesssim$ 100-200 K, with the highest $T_{c}$ seen in hydrides Flores-Livas et al.
+(2020); Duan et al.
+(2017); Pickard et al.
+(2019).
 This naturally stimulates the current intense research into finding materials with $T_{c}$ of 300 K and above.
 We therefore see that the very existence of the current research to find systems with $T_{c}$ above 300 K is itself due to the values of FPCs currently observed.
-3.4 Transitions between states of matter: solids, liquids and gases
+III.4 Transitions between states of matter: solids, liquids and gases
 Three basic states of matter, solids, liquids and gases, are represented by a phase diagram with three first-order phase transition lines as shown in Figure 2.
@@
 At high pressure, melting temperature substantially depends on pressure (see Ref.
-[52] for the functional form of pressure-temperature melting lines).
-At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy [53] because microscopically, melting involves breaking chemical bonds.
+Trachenko (2024) for the functional form of pressure-temperature melting lines).
+At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds.
 This cohesive energy is set by FPCs in Eq.
 (2).
-Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems [53, 54] (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq.
+Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq.
 (12)).
-As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs.
+As discussed in Section III.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs.
 This gives the following role of FPCs in setting the states of matter.
@@
 Conversely, FPCs resulting in small cohesive energies and hence low melting and boiling temperatures imply that matter would exist in the gas state only.
-We see that, similarly to structural phase transitions discussed in Section 3.2, the current values of FPCs set the observability of transitions between different states of matter in Figure 3 and the existence of these states.
-The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life:
+We see that, similarly to structural phase transitions discussed in Section III.2, the current values of FPCs set the observability of transitions between different states of matter in Figure 2 and the existence of these states.
+The apparent overlap of our environmental conditions with characteristic melting and boiling temperatures set by FPCs gives rise to two states important for life:
   1. 1.
@@
 Much smaller values of cohesive energies due to different FPCs would result in H2O and carbon-based structures not existing as cohesive states but as gas states only.
-A more detailed consideration shows an additional and more specific role of FPCs in life processes: FPCs have a biofriendly window constrained by biofriendly viscosity and diffusion setting the motion in essential processes in and across cells [32, 33].
+A more detailed consideration shows an additional and more specific role of FPCs in life processes: FPCs have a biofriendly window constrained by biofriendly viscosity and diffusion setting the motion in essential processes in and across cells Trachenko (2023b, c).
 We note that classical entropies of solids and gases contain FPCs explicitly.
-Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ [46], $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq.
+Recalling the solid entropy $S_{s}\propto 3N\ln\frac{T}{\hbar\omega_{\rm D}}$ Landau and Lifshitz (1970), $\omega_{\rm D}\approx\omega_{m}$ and $\omega_{m}$ in Eq.
 (13) with $m\propto m_{p}$ gives the dependence of $S_{s}$ on FPCs as
 $S_{s}\propto\ln\frac{\hbar^{2}m_{p}^{\frac{1}{2}}}{e^{4}m_{e}^{\frac{3}{2}}}$ (20)
-The entropy of the classical gas is [46]
+The entropy of the classical gas is Landau and Lifshitz (1970)
 $S_{g}=\frac{3}{2}N\left(\ln T+1\right)+N\left(\ln v+\ln\left(\frac{m}{2\pi\hbar^{2}}\right)^{\frac{3}{2}}+1\right)$ (21)
@@
 Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
-Similarly to what we saw in Section 3.2 for structural phase transitions, the emerging high-temperature gas phase increases the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat).
-Similarly to structural transitions, transitions between different states of matter are enabled by FPCs as discussed earlier in this section.
+Similarly to what we saw in Section III.2 for structural phase transitions, the emerging high-temperature gas phase increases the entropy in comparison to what the entropy of the solid would have been in the absence of phase transitions (for first-order transitions, the entropy increase has a contribution from the latent heat).
+Similarly to structural transitions, transitions between different states of matter are constrained FPCs as discussed earlier in this section.
 Differently from solid and gas entropies in Eqs.
 (6), (16) and (21), no equation for the liquid entropy exists due to the fundamental no-small parameter problem of the liquid theory (see Ref.
-[52] for a quick review and Refs.
-[55, 56] for a more detailed discussion).
-Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting [55, 56].
-However, the energy, the primary property in statistical physics [46], can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy $E$ is set by the energies of its excitations [57].
-In liquids, these excitations are gapped phonons and particle jumps [55, 56, 58, 59, 60, 61].
-The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}=\frac{dE}{dT}$ (in almost entire range of liquid relaxation time, $E$ and $c_{v}$ are vibrational to a very good approximation [62]).
+Trachenko (2024) for a quick review and Refs.
+Trachenko and Brazhkin (2016); Trachenko (2023d) for a more detailed discussion).
+Common models such as the hard-sphere and Van der Waals models and their extensions give the constant-volume specific heat $c_{v}=\frac{3}{2}$ and hence describe the ideal gas thermodynamically rather than real liquids where $c_{v}=3$ close to melting Trachenko and Brazhkin (2016); Trachenko (2023d).
+However, the energy, the primary property in statistical physics Landau and Lifshitz (1970), can be calculated in liquids in the closed form using the main premise of statistical physics: the system’s energy $E$ is set by the energies of its excitations Lifshitz and Pitaevskii (2006).
+In liquids, these excitations are gapped phonons and particle jumps Trachenko and Brazhkin (2016); Trachenko (2023d); Proctor (2020, 2021); Chen (2022); de With (2024).
+The entropy can then be found by integrating $\frac{c_{v}}{T}$, where $c_{v}=\frac{dE}{dT}$ (in almost entire range of liquid relaxation time, $E$ and $c_{v}$ are vibrational to a very good approximation Trachenko and Brazhkin (2013)).
 Nevertheless, liquids are a cohesive state, and our discussion of how different states of matter emerge depending on cohesive energies set by FPCs applies to liquids too.
-3.5 Metastable states and chemical reactions
+III.5 Metastable states and chemical reactions
 In the previous sections, we discussed phase transitions and the existence of different states as a result of these transitions.
@@
 Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref.
-[63] for review).
-When discussing transitions in these systems, the tacit assumption is the one we mentioned in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
+Brazhkin (2006) for review).
+When discussing transitions in these systems, the tacit assumption is the one we mentioned in Section III.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
 Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition.
-The world of organic compounds relevant for life processes is nearly all metastable [63].
-Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) [64].
-Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs.
+The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006).
+Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al.
+.
+Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states.
+The kinetics of transitions between metastable states is governed by the bond energies.
+These energies are governed by FPCs as discussed in Section III.
 Hence, similarly to equilibrium transitions, the observability and operation of transitions between different metastable states and the existence of numerous metastable phases is set by FPCs.
@@
 For example, much larger cohesive energy due to different FPCs would imply that chemical reactions and their products (including those essential to life forms) would be inoperable at given environmental parameters.
-4 Fundamental constants and entropy increase
+IV Fundamental constants and entropy increase
 We consider that understanding the current values of FPCs requires the discussion of their variation.
-Theories and observations of this variation were previously discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33].
-Anthropic arguments use this variation to rationalise the observed values of FPCs [9, 10, 13].
-A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) [43] and guided by a unique design and an exquisite symmetry [65].
+Theories and observations of this variation were previously discussed Smolin (1992); Wilczynska et al.
+(2020); Seto et al.
+(2023); Webb et al.
+(2011); King et al.
+(2012); Chiba (2011); Schellekens (2013); Martins (2017); Barrow (2003); Adams (2019); Uzan (2011); Trachenko (2023c).
+Anthropic arguments use this variation to rationalise the observed values of FPCs Carr (2009); Sloan et al.
+(2022); Adams (2019).
+A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007).
 FPCs would then be calculable from this theory.
-In Section 3, we discussed different cases where FPCs, by enabling phase transitions and different phases, increase the entropy.
-We saw that the current values of FPCs enable phase and other transitions, giving rise to different states and the associated increase of the phase space and entropy.
-We also saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy.
+In Section III, we discussed different cases where phase transitions and the emergence of different phases due to FPCs increase the entropy.
+We saw that different values of FPCs would disable these transitions and different phases, resulting in smaller entropy.
 This suggests that the current values of FPCs give rise to entropy increase.
@@
 (17) giving the temperature of phase transitions as a function of FPCs in condensed matter.
-However, the temperature of transitions and reactions depend on FPCs at other energy and length scales as well [6, 13], enabling transitions and reactions at those scales.
+However, the temperature of transitions and reactions depend on FPCs at other energy and length scales as well Barrow (2003); Adams (2019), enabling transitions and reactions at those scales.
 Similarly to condensed matter systems, we expect the new states emerging as a result of these transitions to have higher entropy.
@@
 $d\rho={\rm const}\cdot\delta(E-E_{0})e^{S}\prod\limits_{i}dE_{i}$ (23)
-where $E_{i}$ are energies of subsystems and (b) $d\rho$ is maximised when $E_{i}$ of all subsystems in their own equilibrium states are equal to their most probable equilibrium values $\bar{E_{i}}$, corresponding to full equilibrium [46].
-As discussed in Section 3.2, the subsystem where transitions and reactions take place (phase-changing subsystem, PCS) is not closed and not in equilibrium because these transitions are enabled by the heat flow from another heat-supplying subsystem.
-However, subsystems can be identified which are (a) larger, include both subsystems and are closed and (b) in equilibrium at long post-reaction times [46].
-The conventional mechanism of increase of $S$ and $d\rho$ applies to these closed subsystems, and our entropy increase in the PCS enabled by FPCs additionally increases the statistical probability in Eq.
+where $E_{i}$ are energies of subsystems and (b) $d\rho$ is maximised when $E_{i}$ of all subsystems in their own equilibrium states are equal to their most probable equilibrium values $\bar{E_{i}}$, corresponding to full equilibrium Landau and Lifshitz (1970).
+As discussed in Section III.2, the subsystem where transitions and reactions take place (phase-changing subsystem, PCS) is not closed and not in equilibrium because these transitions are enabled by the heat flow from another heat-supplying subsystem.
+However, subsystems can be identified which are (a) larger, include both subsystems and are closed and (b) in equilibrium at long post-reaction times Landau and Lifshitz (1970).
+The conventional mechanism of increase of $S$ and $d\rho$ applies to these closed subsystems, and our entropy increase in the PCS due to FPCs additionally increases the statistical probability in Eq.
 (23).
@@
 In some processes such as the above example of nuclear synthesis in stars, the environmental parameters can also change as a result of reactions themselves.
-Nevertheless, the transitions and reactions are enabled or disabled by FPCs at varying environmental parameters as they are at fixed parameters.
-We recall the discussion of long-lived metastable states with higher entropy in Section 3.
-As recently reviewed [66], metastable states are ubiquitous at different levels of matter organization in the Universe including nuclei and astronomical objects.
-The energy barriers setting the lifetime of these states vary in a wide range 10-7 to 109 eV per particle, with relaxation times often exceeding the age of the Universe [66].
-These metastable states, similarly to those discussed in Section 3, emerge as a result of different transitions and reactions which depend on the values of FPCs [6, 13] and can co-exist at the same environmental parameters.
-The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution due to environmental pressure.
-Based on this entropy increase and the associated increase of statistical probability (22), we conjecture that entropy increase is a selection principle for FPCs considered to vary in earlier discussions (see, e.g., Refs.
-[35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 4, 33] and references therein).
+Nevertheless, the transitions and reactions are constrained by FPCs at varying environmental parameters as they are at fixed parameters.
+We recall the discussion of long-lived metastable states with higher entropy in Section III.
+As recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels of matter organization in the Universe including nuclei and astronomical objects.
+The energy barriers setting the lifetime of these states vary in a wide range 10-7 to 109 eV per particle, with relaxation times often exceeding the age of the Universe Brazhkin (2024).
+These metastable states, similarly to those discussed in Section III, emerge as a result of different transitions and reactions which depend on the values of FPCs Barrow (2003); Adams (2019) and can co-exist at the same environmental parameters.
+The entropy increase due to FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution due to environmental pressure.
+Based on this entropy increase and the associated increase of statistical probability (23), we conjecture that total entropy increase is a selection principle for FPCs considered to vary in earlier discussions (see, e.g., Refs.
+Smolin (1992); Wilczynska et al.
+(2020); Seto et al.
+(2023); Webb et al.
+(2011); King et al.
+(2012); Chiba (2011); Schellekens (2013); Martins (2017); Barrow (2003); Adams (2019); Uzan (2011); Trachenko (2023c) and references therein).
 In other words, FPCs tend to give the most probable state.
-The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, is unknown, although several proposals were discussed [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33].
+The mechanism of variability or fluctuations of FPCs, analogous to genetic mutation in biological evolution, is unknown, although several proposals were discussed Smolin (1992); Wilczynska et al.
+(2020); Seto et al.
+(2023); Webb et al.
+(2011); King et al.
+(2012); Chiba (2011); Schellekens (2013); Martins (2017); Barrow (2003); Adams (2019); Trachenko (2023c).
 The above conjecture is nevertheless interesting to consider because, if correct, it is analogous to introducing a Darwinian concept of a selection principle, the notion that brought new understanding even though the molecular DNA-based mutation mechanism was not known at the time and came about a century later only.
 Cyclic or oscillatory models of the Universe (see, e.g., Refs.
-[67, 35, 68, 69, 70, 71, 72] and Ref.
-[73] for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle.
+Adams and Laughlin (1997); Smolin (1992); Steinhardt and Turok (2002); Baum and Frampton (2007); Ashtekar and Singh (2011); Gurzadyan and Penrose (2013); Trachenko (2017) and Ref.
+Kragh (2009) for a review including epistemic and philosophical aspects) can account for inheriting FPCs from one cycle to the next and maintain fluctuations around their values in the previous cycle.
 In these models, the change of FPCs is often related to the smallest size of the Universe and very high energy including the Planck energy scale where unknown physics operates.
 The size of the oscillating Universe at the inversion point in Ref.
-[72] is set by the system energy as a parameter in the model and does not need to be small.
+Trachenko (2017) is set by the system energy as a parameter in the model and does not need to be small.
 We might ask whether the current FPCs give structures in the Universe that correspond to the maximised phase space and entropy or close to it.
-Adams concludes [13] that the observable Universe is not the most abundant and complex possible: somewhat different values of FPCs would give more complex and abundant structures (e.g.
+Adams concludes Adams (2019) that the observable Universe is not the most abundant and complex possible: somewhat different values of FPCs would give more complex and abundant structures (e.g.
 more baryons and stars, more efficient nuclear synthesis and so on).
 Another interesting observation in Ref.
-[13] is that our Universe does not lie at the centre of allowed range of values of FPCs.
+Adams (2019) is that our Universe does not lie at the centre of allowed range of values of FPCs.
 The evolutionary mechanism with the entropy increase as a selection principle has several potential implications for these observations.
-We first note that this selection applies to FPCs and FPC-enabled structures existing at the time (as is the case in biological evolution) rather than to currently observed structures.
-Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist [13], resulting in scarcity of structures and smaller phase space.
+We first note that this selection applies to FPCs and structures appearing due to these FPCs existing at the time (as is the case in biological evolution) rather than to currently observed structures.
+Second, the observed structures depend on the magnitude of fluctuations of FPCs: large fluctuations destroy the intricate network of constraints on the values of FPCs needed for different structures to exist Adams (2019), resulting in scarcity of structures and smaller phase space.
 On the other hand, smaller fluctuations maintain the continuity of structures in different cycles.
@@
 Third, it follows that maximising entropy as a function of FPCs is subject to many constraints on their values.
-This invites an input from the optimisation theory.
+Hence understanding the values of FPCs (e.g.
+calculating these values or their ranges) invites an input from the optimisation theory.
+This programme would include taking an initial set of FPCs as variables, calculating the entropies of important structures and formations in the Universe and finding FPCs or their ranges by maximising the total entropy while (a) accounting for constraints on FPCs from each structure Barrow (2003); Adams (2019) and (b) remembering that later-appearing structures may rely on the existence of earlier ones.
+This process prevents, for example, finding FPCs giving larger entropy of earlier structures but smaller entropy of later structures (or no entropy of later structures if they are not formed due to unfavourable FPCs involved in the emergence of earlier structures) or FPCs giving larger entropy of later structures but neglecting earlier structures needed for later structures to emerge.
 An example related to this discussion and involving constraints is the black hole (BH) entropy.
-If most of the current entropy in the Universe is associated with the entropy of large BHs [74], one might consider small $\hbar$ and $c$ to maximise the BH entropy
+If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
 $S_{BH}=4\pi\frac{G}{\hbar c}M_{BH}^{2}$ (24)
@@
 However, small $\hbar$ and $c$ give large $\alpha$ in Eq.
-(2), breaking many constraints required for key processes of structure formation and disrupting those processes [6, 13].
-This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times [6, 13] (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
-We also note that considering variation of the BH entropy due to different FPCs is unimportant for the evolution of FPCs at earlier stages when BHs may not have existed yet: as mentioned earlier, this evolution applies to structures existing at the time only.
+(4), breaking many constraints required for key processes of structure formation and disrupting those processes Barrow (2003); Adams (2019).
+This includes processes preceding the formation of BHs such as stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
 The discussion of FPCs acting to increase the entropy and the conjecture of entropy increase as a selection principle for FPCs need further study.
-This includes detailing entropy changes in processes at different levels of structure organisation in the Universe [6, 13], identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity [46] and so on.
-Given the scale of the problem we face in understanding the origin of fundamental constants [6, 15, 14], making new observations and asking new questions is useful for progress.
-5 Discussion and summary
+This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed or conditionally closed and where the entropy argument applies, clarifying the role of gravity Landau and Lifshitz (1970) and so on.
+Given the scale of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.
+V Discussion and summary
 Approaching the challenge to explain fundamental constants started with understanding the role these constants play.
-This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures [6, 7, 8, 9, 10, 11, 12, 13, 3].
+This role has historically been limited to high-energy nuclear physics and astrophysics physics and corresponding structures Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al.
+(2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003).
 Condensed matter physics was relatively unexplored in this regard, however more recently it was realised that FPCs set lower and upper bounds on key condensed matter properties.
@@
 Ref.
-[31] (pages 479-480) lists important differences between such a theory-based approach and a dimensional analysis.
+Trachenko (2023a) (pages 479-480) lists important differences between such a theory-based approach and a dimensional analysis.
 We saw that were FPCs to take different values, transitions would be disabled at our environmental conditions and new different phases would not emerge.
@@
 In this sense, the insights into FPCs from condensed matter physics face less uncertainty and challenges as compared to models in high-energy physics and astrophysics (see Ref.
-[13] for a review of these challenges).
+Adams (2019) for a review of these challenges).
 If the insights into the role of FPCs in condensed matter physics are found to be general enough, these insights can be applied to structures at other levels in the Universe.
@@
 Last but not least, deeper understanding the role of FPCs in condensed matter physics is important on its own.
-A fairly recent realisation that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models.
-These properties include viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34].
+A realization that FPCs put lower or upper bounds on key condensed matter properties has deepened our understanding of how these properties operate and benefited our physical models.
+These properties include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al.
+(2017); Sommer et al.
+(2011); Bardon et al.
+(2014); Trotzky et al.
+(2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022); Grozdanov (2021), thermal conductivity in insulators Trachenko et al.
+(2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al.
+(2020), elastic moduli Brazhkin (2023); Trachenko (2023a), flow in cells Trachenko (2023b, c), vibration frequency in solids and superconducting temperature Trachenko et al.
+.
 Understanding how FPCs govern the observability and operation of entirely new effects and phenomena is another useful insight enhancing our understanding of condensed matter physics.